### v1

# CLIMB-3D: Continual Learning for Imbalanced 3D Instance Segmentation

Abstract
While 3D instance segmentation has made significant progress, current methods struggle to address realistic scenarios where new categories emerge over time with natural class imbalance. This limitation stems from existing datasets, which typically feature few well-balanced classes. Although few datasets include unbalanced class annotations, they lack the diverse incremental scenarios necessary for evaluating methods under incremental settings. Addressing these challenges requires frameworks that handle both incremental learning and class imbalance. However, existing methods for 3D incremental segmentation rely heavily on large exemplar replay, focusing only on incremental learning while neglecting class imbalance. Moreover, frequency-based tuning for balanced learning is impractical in these setups due to the lack of prior class statistics. To overcome these limitations, we propose a framework to tackle both Continual Learning and class Imbalance for 3D instance segmentation (CLIMB-3D). Our proposed approach combines Exemplar Replay (ER), Knowledge Distillation (KD), and a novel Imbalance Correction (IC) module. Unlike prior methods, our framework minimizes ER usage, with KD preventing forgetting and supporting the IC module in compiling past class statistics to balance learning of rare classes during incremental updates. To evaluate our framework, we design three incremental scenarios based on class frequency, semantic similarity, and random grouping that aim to mirror real-world dynamics in 3D environments. Experimental results show that our proposed framework achieves state-of-the-art performance, with an increase of up to 16.76% in mAP compared to the baseline. Code is available at: https://github.com/vgthengane/CLIMB3D
1 Introduction
Figure 1: Class-incremental setup for 3D instance segmentation. As tasks progress through time, new classes are introduced incrementally. After each new task, the model should recognize both previously learned and newly introduced classes. For example, at Task-2, new classes such as Pillow, Coffee Table, and Sofa Chair are added, and the model is able to detect these three classes along with previous ones like Table, Chair, and Couch.
3D instance segmentation is an essential task in computer vision that involves identifying and segmenting individual objects in the real physical space, playing a key role in applications across graphics, robotics, and autonomous systems. Its ability to provide precise object boundaries and class labels enhances scene understanding, facilitates object manipulation, and improves perception in dynamic environments.
In recent years, a variety of methods have been proposed, including top-down approaches [wang2018sgpn, jia2021scaling, zhang2021point], bottom-up approaches [hou20193dsis, yang2019learning], and transformer-based architectures [Schult23ICRA]. These methods have shown impressive results in traditional setups, which assume that all object classes are available during training. However, this assumption limits its applicability in real-world scenarios where new categories gradually emerge over time, often exhibiting naturally imbalanced distributions. Thus, there is a need for class-incremental learning (CIL) frameworks that not only adapt to new classes but also preserve prior knowledge, especially for rare or less frequent categories, which are more prone to catastrophic forgetting.
Most existing research in class-incremental learning focuses on 2D image classification [rebuffi2017icarl, li2017learning, aljundi2018memory, serra2018overcoming], with some extensions to object detection [liu2023continual, joseph2021towards, shmelkov2017incremental] and semantic segmentation [cermelli2020modeling, cermelli2022incremental, douillard2021plop]. These methods employ strategies such as exemplar replay [buzzega2020dark, rebuffi2017icarl, cha2021co2l, kamra2017deep], regularization [aljundi2018memory, li2017learning, serra2018overcoming], and knowledge distillation [douillard2020podnet, kang2022class] to preserve previously learned knowledge and mitigate catastrophic forgetting [mccloskey1989catastrophic]. Few studies have applied CIL to point clouds; however, they mostly focus on object-level classification [dong2021i3dol, liu2021l3doc, chowdhury2021learning]. At the scene level, some works have explored 3D semantic segmentation [Yang_2023_CVPR] with incremental learning, but their performance is not as competitive as state-of-the-art methods that do not employ incremental learning, which limits their applicability. Other methods tackle open-world incremental learning [boudjoghra20243d] but rely heavily on large exemplar sets [rolnick2019experience] and often neglect the challenge of class imbalance.
To address this, we propose CLIMB-3D, a unified framework that combines Exemplar Replay (ER), Knowledge Distillation (KD), and a novel Imbalance Correction (IC) module to tackle Continual Learning for Imbalance 3D instance segmentation in indoor environments, as shown in Figure 1. Our framework operates as follows: ER stores a subset of representative samples from previous stages, allowing for effective replay during new task learning. KD transfers knowledge by retaining a copy of the model from the previous task, thereby mitigating forgetting. The IC module is specifically designed to reduce forgetting of rare classes by leveraging the frequency of object occurrence. However, since we do not have access to the previous task data or statistics during incremental phases, we instead compile these statistics from the previous model used by KD to generate weight for earlier categories. These weights are used to favor both frequent and rare classes, ensuring a balanced learning and mitigating forgetting.
To evaluate CLIMB-3D in a realistic, incremental learning setup, we create three benchmark scenarios using the ScanNet200 dataset [rozenberszki2022language], which features 200 classes with natural class imbalances. These scenarios are designed to reflect real-world conditions where new categories emerge gradually and follow inherent class imbalances. These are based on 1 frequency of object occurrence, 2 semantic similarity between the object, and 3 random grouping. Our experiments demonstrate that CLIMB-3D significantly improves performance by effectively mitigating the forgetting of previous tasks compared to earlier exploration in class incremental 3D segmentation.
In summary, our contributions are as follows:
  1. 1.
    We propose a new problem setting for imbalanced class incremental 3D segmentation, along with a simple yet effective method to address this challenge by minimizing catastrophic forgetting and balancing the learning process.
  2. 2.
    To benchmark this setting, we design three scenarios aimed at simulating real-world conditions where objects emerge continuously with natural class imbalance.
  3. 3.
    Experimental results show that our proposed framework achieves state-of-the-art performance, with an increase of up to 16.76% in mAP compared to the baseline.
2 Related Work
This section reviews the current literature on 3D instance segmentation and incremental learning methods, including the limited work addressing incremental learning for 3D scene-level tasks.
2.1 3D Instance Segmentation
Various approaches have been proposed for 3D instance segmentation. One common approach adopts a bottom-up pipeline, in which an embedding in the latent space is learned to facilitate the clustering of object points [wang2018sgpn, lahoud20193d, jiang2020pointgroup, zhang2021point, chen2021hierarchical, han2020occuseg, he2021dyco3d, liang2021instance]. These methods are also known as grouping-based or clustering-based methods. Other methods use a top-down approach, also known as proposal-based methods, where 3D bounding boxes are first detected, then the object region is segmented within the box [yang2019learning, hou20193dsis, liu2020learning, yi2019gspn, engelmann20203d]. Recently, the transformer architecture [vaswani2017attention] has also been used for the task of 3D instance segmentation [Schult23ICRA, sun2022superpoint], motivated by work in 2D [cheng2022masked, cheng2021per]. While these methods propose various models for improving the quality of the object segments, they rely on the availability of annotations for all object categories. On the other hand, we target learning in a progressive manner, in which new semantic annotation is provided and past data is inaccessible.
In order to reduce the annotation cost for 3D instance segmentation, various methods propose weakly supervised alternatives to methods that use dense annotations [xie2020pointcontrast, hou2021exploring, chibane2022box2mask]. While these methods improve the ability to learn from a small set of annotated examples, they rely on a fixed set of semantic labels, so they are prone to catastrophic forgetting in an incremental setting.
2.2 Incremental Learning
Incremental, lifelong, or continual learning methods aim to train a machine learning model sequentially to avoid “catastrophic forgetting” which is caused by training the model on a set of data and later training on another set of data. There are several methods have been proposed for this paradigm, these methods can be divided into three categories: (i) Model Regularization [kirkpatrick2017overcoming, aljundi2018memory, li2017learning] methods limit the plasticity of model parameters to avoid catastrophic forgetting of previous tasks. These methods include weight regularization such as EWC [serra2018overcoming] and function regularization such as knowledge distillation [hinton2015distilling]. (ii) Exemplar replay approaches either create a subset of the past task data or generate samples using generative models to avoid privacy concerns and save those in memory to replay while learning new tasks. This method is effective in more challenging settings and datasets [rebuffi2017icarl, kamra2017deep, buzzega2020dark, cha2021co2l]. (iii) Dynamic network expansion-based method learns a new task by either dynamically expanding the model [rusu2016progressive, li2019learn, zhao2022deep] or by creating a subset of the model [zhao2022deep, wang2020learn, ke2020continual, rajasegaran2019random] to learn to cater for a new task.
Recent approaches to 3D class-incremental segmentation, such as [Yang_2023_CVPR] and [su2024balanced], have made some initial contributions. However, these methods often fall short in performance as they do not leverage state-of-the-art 3D segmentation models and are primarily focused on semantic segmentation, while our work emphasizes object-level instance segmentation. Kontogianni et al. [kontogianni2024continual] propose a general online continual learning framework and evaluate it on 3D dataset segmentation. Similarly, [boudjoghra20243d] addresses the open-world 3D incremental learning problem but relies heavily on an extensive memory buffer. In contrast, our work introduces a dedicated continual learning framework for 3D instance segmentation, with a focus on effective knowledge transfer from previous tasks, while also accounting for the challenges posed by infrequent class occurrences.
3 Preliminaries
3.1 Transformer-based Segmentation
We adopt a transformer-based instance segmentation method based on Mask3D [Schult23ICRA]. Specifically, transformer-based segmentation model $\Phi$ is employed for point cloud instance segmentation. Given an input point cloud $p$, the model predicts $\hat{y}=\{(\hat{m_{j}},\hat{c_{j}})\}_{j=1}^{J}$, which consists of mask predictions and class probabilities for each instance. The segmentation process begins by quantizing the input point cloud $p$ into voxels $V$, creating voxelized representations of size $\mathbb{R}^{M_{0}\times 3}$. Each voxel is assigned an average RGB color computed from the points within that voxel, serving as its initial feature representation. The feature backbone network generates a high-resolution output feature map $\mathbf{F}_{0}\in\mathbb{R}^{M_{0}}$. Additionally, intermediate feature maps are extracted from the decoder layers of the backbone network. For each intermediate feature map ($r\geq 0$), a set of $K_{r}$ voxels is selected, and their features are linearly projected to a fixed dimension $D$, yielding feature maps $\mathbf{F}_{r}\in\mathbb{R}^{M_{r}\times D}$.
The Transformer decoder initiates with a set of K instance queries and iteratively improves them using L Transformer decoder layers. These layers employ cross-attention to refine the instance queries, incorporating information from point cloud features. The decoder attends to a specific feature map obtained from the corresponding feature backbone layer at each layer, employing conventional cross-attention mechanisms. This process enables the decoder to reason at the instance level through self-attention, resulting in the generation of accurate and contextually relevant instance queries tailored to the specific scene.
To achieve this, the voxel features $\mathbf{F}_{r}\in\mathbb{R}^{M_{r}\times D}$ are transformed into sets of keys $\mathbf{K}\in\mathbb{R}^{M_{r}\times D}$ and values $\mathbf{V}\in\mathbb{R}^{M_{r}\times D}$ through linear projection. The instance queries $\mathbf{Z}$ are also projected to create the queries $\mathbf{Q}$. This enables cross-attention, allowing the queries to gather relevant information from the voxel features. Following cross-attention, a self-attention step occurs among the queries, facilitating information exchange and refinement. The learned queries are then used to make $K$ class and mask predictions, which are matched with ground truth labels through bipartite matching, resulting in $\hat{y}=\{(\hat{m}_{j},\hat{c}_{j})\}_{j=1}^{J}$. The model is optimized based on the ground truth label, mask, and class predictions:
$\mathcal{L}_{\mathrm{Seg}}(y_{j},\hat{y}_{j})=\mathcal{L}_{\mathrm{mask}}(m_{j},\hat{m}_{j})+\lambda_{\mathrm{cls}}\mathcal{L}_{\mathrm{cls}}(c_{j},\hat{c}_{j})$ (1)
where, mask loss $\mathcal{L}_{\mathrm{mask}}=\lambda_{\mathrm{ce}}\mathcal{L}_{\mathrm{ce}}(y_{j},\hat{y}_{j})+\lambda_{\mathrm{dice}}\mathcal{L}_{\mathrm{dice}}(y_{j},\hat{y}_{j})$ and $\lambda_{\mathrm{cls}}\mathcal{L}_{\mathrm{cls}}$ is classification loss.
The traditional setup assumes all categories are available and well-balanced during training. However, in scenarios where only a subset of categories is present, training the model in multiple phases is required. Unfortunately, such multi-phase models often suffer from forgetting previous tasks. To address this issue, we employ incremental learning strategies which will be discussed in the next section.
4 Methodology
In this section, we begin by formulating the incremental 3D instance segmentation problem and introduce our proposed method to address it using a transformer-based model. We then detail the design of incremental scenarios, outlining the motivation and considerations behind each one.
Figure 2: CLIMB-3D: At phase $t$ (for $t>1$), we create a copy of the model from the previous phase, $\Phi^{\mathrm{T}}$ (the teacher model), and designate it as $\Phi^{\mathrm{S}}$ (the student model). Both $\Phi^{\mathrm{T}}$ and $\Phi^{\mathrm{S}}$ process the input point cloud simultaneously, producing predictions $\hat{y}_{\mathrm{pseudo}}$ and $\hat{y}$, respectively. To prevent the model from forgetting less frequent categories from previous tasks, we balance the predictions from $\Phi^{\mathrm{T}}$, resulting in a balanced output $\hat{y}^{\mathrm{(w)}}_{\mathrm{pseudo}}$. The concatenated vector of ground truth labels $[y,\hat{y}_{\mathrm{pseudo}},\hat{y}^{\mathrm{(w)}}_{\mathrm{pseudo}}]$ is then compared with the predicted labels $\hat{y}$. A loss function is applied based on this comparison, enabling the student model to learn from the differences between the predicted and the concatenated pseudo-ground truth labels
4.1 Problem Formulation
The objective of 3D point cloud instance segmentation is to accurately identify and segment individual instances of objects within a given point cloud. Mathematically, the training dataset is represented as $\mathcal{D}=(\mathcal{P},\mathcal{Y})=\{(p_{i},y_{i})\}_{i=1}^{N}$, where $N$ is the total number of samples. Each sample consists of a colored point cloud $p_{i}\in\mathbb{R}^{M\times 6}$ of size $M$, where the point coordinates and color values are represented as $\{x,y,z,r,g,b\}$. The corresponding annotations are denoted as $y_{i}=\{(m_{i,j},c_{i,j})\}_{j=1}^{J}$, where $m_{i,j}$ represents the instance mask for the $j$-th instance, and $c_{i,j}\in\mathcal{C}=\{1,\dots,C\}$ denotes the semantic label of the object category to which the instance belongs for the $i$-th point cloud. Here, $J$ represents the total number of instances in the $i$-th point cloud, and $C$ indicates the number of distinct object categories. During the learning process, the model $\Phi$ will process this dataset and output predictions $\hat{y}_{i,j}=(\hat{m}_{i,j},\hat{c}_{i,j})$, where $\hat{m}_{i,j}$ represents the predicted instance mask and $\hat{c}_{i,j}$ denotes the predicted semantic label for the $j$-th instance in the $i$-th point cloud.
To adapt the dataset to an incremental learning setting, we partition the object categories $\mathcal{C}$ into $T$ subsets, denoted as $\mathcal{C}=\mathcal{C}^{1}\cup\cdots\cup\mathcal{C}^{T}$. Each phase $t\in\{1,\dots,T\}$ is associated with a specific subset $\mathcal{C}^{t}$, and its corresponding dataset $\mathcal{D}^{t}$ which only contains annotations for objects belonging to the corresponding subset. Formally, during the $t$-th phase of training, the dataset $\mathcal{D}^{t}=(\mathcal{P},\mathcal{Y}^{t})=\{(p_{i},y_{i}^{t})\}_{i=1}^{N}$ is defined, where $\mathcal{P}$ represents the point clouds shared across all phases, and $\mathcal{Y}^{t}$ contains annotations exclusively for objects belonging to the class subset $\mathcal{C}^{t}$. It is important to note that the 3D scenes within each phase can contain objects of any type from the entire object category set $\mathcal{C}$, but only the object belonging to $\mathcal{C}^{t}$ are annotated during that specific phase. After training for phase $t$ completes, the model is evaluated on a validation set containing the union of classes up to task $t$ (i.e., $\mathcal{C}^{1}\cup\cdots\cup\mathcal{C}^{t}$). Training progresses to the next phase, $t+1$, where the model $\Phi$ observes the same set of 3D scenes $\mathcal{P}$ but with annotations for different object types belonging to the subset $\mathcal{C}^{t+1}$. This incremental training approach allows the model to gradually learn and adapt to new object categories over multiple phases.
4.2 CLIMB-3D
In our proposed framework (Figure 2), the incremental instance segmentation model undergoes phased training, as described in Sec. 4.1, where carefully designed subsets of the dataset are introduced to handle various real-world scenarios discussed in Sec. 4.3. Formally, at phase $t$, when the model $\Phi^{t}$ is introduced with input data $\mathcal{D}^{t}=\{(p_{i},y_{i}^{t})\}_{i=1}^{N}$ and trained using Eq. 1, a common issue arises where it tends to forget the knowledge acquired in the previous phase, leading to catastrophic forgetting [mccloskey1989catastrophic]. To address this, we first draw inspiration from techniques developed in the 2D domain [rebuffi2017icarl, li2017learning] and recent 3D semantic segmentation [su2024balanced, Yang_2023_CVPR], and adapt them for our setting. However, we observe that these adaptations alone fall short of achieving the desired performance levels; therefore, we propose a teacher-student knowledge distillation framework to effectively retain previously learned knowledge. Additionally, we incorporate an imbalance correction module to handle the challenge of less frequent classes from earlier tasks.
Exemplar Replay (ER). Inspired by the approach proposed by Buzzega et al. [buzzega2020dark], ER methods alleviate the issue of limited exposure to previous task data during training. By selectively storing a small subset of exemplars $\mathcal{E}_{t}$ from previous phases, the model can learn from both the current task data $\mathcal{D}_{t}$ and the replayed exemplars $\mathcal{E}_{1:t-1}$. This results in a combined dataset $\mathcal{D}_{t}\cup\mathcal{E}_{1:t-1}$, where $\mathcal{E}_{1:t-1}$ represents the exemplar memory formed by the union of all previous exemplar sets $\mathcal{E}_{1:t-1}=\mathcal{E}_{1}\cup~{}\dots~{}\cup\mathcal{E}_{t-1}$. The model undergoes a full iteration on $\mathcal{D}_{t}$ before replaying the exemplars.
Some previous 3D approaches adopt this strategy to retain knowledge but rely on a large exemplar set [boudjoghra20243d], which is often impractical in real-world scenarios. To address this, we choose a smaller exemplar set, creating a more challenging setup that requires the model to effectively manage and retain knowledge with limited resources, thereby testing its robustness and adaptability in practical applications.
Knowledge Distillation (KD) Module. In our incremental learning approach, we utilize a Knowledge Distillation (KD) module that incorporates a teacher-student framework, maintaining a copy of the previously trained model while learning the current task. For $t>1$, at the beginning of each training stage, the current model $\Phi^{t}$ is initialized as $\Phi^{t}\leftarrow\Phi^{t-1}$, where $\Phi^{t-1}$ represents the model trained in the previous phase. As the $\Phi^{t}$ is trained on the previous stage dataset, it holds information about the previous set of classes. Hence, we make use of this model to help retain previous knowledge while learning the current task.
When presented with a new training point-cloud and label pair $(p,y^{t})$, the output of the previous model is calculated as $\hat{y}_{\mathrm{pseudo}}=\Phi^{\mathrm{t-1}}(p)$, and a combined loss function is minimized. This combined loss comprises the $\mathcal{L}_{\mathrm{Seg}}(\hat{y}^{t},y^{t})$ loss, which measures the discrepancy between the predicted and ground truth labels, and the knowledge distillation loss $\mathcal{L}_{\mathrm{KD}}$, which encourages the similarity between the predictions of the current model and the previous model.
$\mathcal{L}_{\mathrm{KD}}(\hat{y}^{t},\hat{y}_{\mathrm{pseudo}})$ $=\mathcal{L}_{\mathrm{mask}}(\hat{m}^{t}_{j},\hat{m}_{j,\mathrm{pseudo}})$
$\quad\quad\quad\quad+\lambda_{\mathrm{cls}}\mathcal{L}_{\mathrm{cls}}(\hat{c}^{t}_{j},\hat{c}_{j,\mathrm{pseudo}})$ (2)
However, as pointed out in previous works on object detection in 2D [liu2023continual], Section 4.2 is often biased towards the background classes, as the model tends to predict the background for most instances. Similarly, we propose selecting the top $K$ most confident predictions from the previous teacher model $\Phi^{t-1}$ and combining them with the ground truth labels, which then serve as pseudo-labels. By extracting the top $K$ confident samples from the output of $\Phi^{t-1}$ and combining them with the ground truth labels, the augmented label set becomes $y^{\prime}=[y^{t},\hat{y}_{\mathrm{pseudo}}^{K}]$. This augmented label set is then used to optimize the current model $\Phi$ with the $\mathcal{L}_{\mathrm{Seg}}$ loss function from Equation 1.
Imbalance Correction (IC) Module. Although retaining a few samples from previous tasks and selecting the most confident predictions from the previous model, $\Phi^{t-1}$, helps preserve information from prior tasks while learning new ones, we observed that this approach does not adequately address the class imbalance. Our analysis reveals that the most confident predictions from $\Phi^{t-1}$ are largely associated with the most frequent object categories, causing the model to forget less common classes. This issue can be mitigated by re-weighting the predictions based on the frequency of observed categories [islam2021class].
At task $t$, we only have access to the data and statistics of the current task; the previous task’s dataset and statistics are unavailable. To incorporate balancing elements despite the absence of previous stage statistics, we propose leveraging pseudo prediction statistics from $\Phi^{t-1}$. During each iteration, we use $\Phi^{t-1}$ to generate pseudo labels and accumulate class frequency statistics for prior tasks throughout the current epoch training. At the end of each epoch, we combine the statistics of observed scene classes and predicted pseudo-classes, calculating the frequency $\mathbf{f}$ of all classes seen so far.
Formally, for each category $c$, we assign a weight $\mathbf{w}_{c}$ inversely proportional to its observed frequency in the predictions of $\Phi^{t-1}$ and the current dataset. The weight $\mathbf{w}_{c}$ is defined as: $\mathbf{w}_{c}=\frac{1}{\mathbf{f}(c)+\epsilon}$, where $\epsilon$ is a small constant to avoid division by zero. In the next epoch, predictions from $\Phi^{t-1}$ are re-weighted using $\mathbf{w}_{c}$, creating an adjusted high-confidence pseudo label set for less frequent categories: $\hat{y}_{\mathrm{pseudo}}^{(w)}=\mathbf{w}_{c}\cdot\hat{y}_{\mathrm{pseudo}}$. This re-weighting occurs at each epoch, allowing $\Phi^{t-1}$ to yield a broader set of less frequent classes. To ensure the model encounters both high-confidence and less frequent classes, we select the top $K$ high-confidence predictions both before and after re-weighting. The resulting augmented label space, which combines ground truth labels, original pseudo labels, and re-weighted pseudo labels, is given by: $y^{\prime}=[y,\hat{y}_{\mathrm{pseudo}}^{K},y_{\mathrm{pseudo}}^{(w),K}]$.
As this weighting scheme is applied only to previous model predictions, we further tune the current model to favor less frequent classes by incorporating the same weights $\mathbf{w}_{c}$ into the classification loss of Equation 1. The adjusted segmentation loss becomes:
$\mathcal{L}_{\mathrm{3DIS}}(y^{\prime}_{j},\hat{y}_{j})=\mathcal{L}_{\mathrm{mask}}(m^{\prime}_{j},\hat{m}_{j})+\mathbf{w}^{\prime}_{c}\mathcal{L}_{\mathrm{cls}}(c^{\prime}_{j},\hat{c}_{j}),$ (3)
where $\mathbf{w}^{\prime}_{c}=\mathbf{w}_{c}\cdot\lambda_{\mathrm{cls}}$ represents the adjusted category weights. By re-weighting and augmenting both the label space and the loss function, our IC module ensures that both the current and previous models encounter pseudo labels spanning the long-tail distribution, addressing class imbalance and enabling more balanced learning.
4.3 Designing Incremental Scenarios
Figure 3: Incremental scenarios are grouped based on frequency of occurrence, semantic similarity, and random clustering. Different color clouds ( ,  ,  ) represent tasks in each scenario, while various shapes represent object categories, and   denotes the background. Left: Tasks are organized based on the frequency of object categories. Middle: Tasks are grouped by semantic similarity, where objects with similar shapes (e.g., circles, plus signs, and triangles) denote semantically similar classes. Right: In this fully random scenario, tasks may contain a mix of semantically similar, more frequent, or less frequent classes.
While conventional incremental learning methods have numerous practical applications, they often assume an equal distribution of samples, which does not reflect real-world conditions. In practice, the number of object categories, $\mathcal{C}$, is typically large, with significant variability in category occurrence, shape, structure, and size. With these attributes in mind, we design three incremental learning scenarios, each addressing distinct aspects of real-world conditions, the design is highlighted in Figure 3.
1
Frequency Scenarios (Split_A). This scenario acknowledges that datasets are often labeled based on the frequency of category occurrences. To accommodate this, we propose a split where the model initially learns from the most frequent categories and subsequently incorporates the less frequent ones in later stages. By prioritizing the training of frequently occurring categories, the model can establish a strong foundation before expanding its knowledge to handle rarer categories.
2
Semantic Scenarios (Split_B). In real-world environments, objects may exhibit similarities in appearance, and then moved to different environments, the model may encounter new objects that does not share similar semantic characteristics with previously seen categories. To address this, we introduce the Split_B  scenario. It involves grouping categories based on their semantic labels and incrementally training the model to handle these groups. This allows the model to generalize its knowledge across semantically similar categories, facilitating adaptation to new objects with similar characteristics. Unlike the Split_A  scenario, this scenario may include both frequent and infrequent categories within the same task.
3
Random Scenarios (Split_C). In some cases, data labeling is based on the availability of objects rather than specific criteria. To account for this scenario, we design the Split_C  scenario. This scenario represents a completely random setting where each task can have any class, leading to varying degrees of class imbalance. By exposing the model to such diverse and imbalanced scenarios, we aim to enhance its ability to handle real-world situations where the availability of labeled data is unpredictable.
By designing these three incremental learning scenarios, we aim to provide a more realistic representation of object distributions, frequencies, and dynamics encountered in the real world.
5 Experiments
We evaluate our method across three scenarios specifically designed to capture the complexities of real-world settings, where new categories emerge incrementally alongside class imbalance, as discussed in Section 4.3. Experimental results demonstrate that our approach effectively handles the gradual introduction of new classes and mitigates the impact of class imbalance within these scenarios. The following sections detail the datasets, evaluation metrics, incremental scenarios, and implementation procedures, followed by a thorough analysis and comparison of the results.
5.1 Experimental Setup
Datasets. We evaluate our method using the ScanNet200 dataset [rozenberszki2022language], which includes 200 object categories and exhibits inherent class imbalance, making it well-suited for simulating and evaluating real-world scenarios. Additionally, we benchmark our method against existing incremental learning approaches using the original ScanNet dataset [dai2017scannet] in a semantic segmentation setting. We follow the standard training and validation splits as defined in prior works.
Evaluation Metrics. We evaluate our method using mean Average Precision (mAP), a standard metric for 3D instance segmentation that provides a comprehensive measure of segmentation quality, accounting for both precision and recall. For comparison with existing semantic incremental learning approaches, we report mean Intersection over Union (mIoU), which measures the overlap between predicted and ground truth instances, offering a detailed evaluation of segmentation accuracy. To assess the model’s ability to mitigate catastrophic forgetting in continual learning scenarios, we use the Forgetting Percentage Points (FPP), as defined in [liu2023continual]. This metric quantifies performance degradation by measuring the accuracy drop between the initial and final training phases, on the categories observed in the first training phase.
Incremental Scenarios. As discussed in Section 4.3, we design three incremental scenarios: Split_A, Split_B, and Split_C, each consisting of three tasks, which are grouped based on object occurrence frequency, semantic similarity, and random grouping, respectively. In Split_A, the frequency of object categories progressively decreases through the tasks. This scenario follows the head, common, and tail splits present in the ScanNet200 dataset, with class distributions of 66-68-66 in each split. In Split_B, we partition the classes into 74-50-76 based on semantic similarity, which is calculated using the CLIP [radford2021learning] text encoder, followed by clustering using K-Means. Finally, in Split_C, the classes are shuffled and split into three sets, resulting in 67-67-66 categories per split. These scenarios allow for a comprehensive evaluation of our approach under varying conditions, facilitating a deeper understanding of its performance and generalization in diverse real-world settings.
Implementation Details. We utilize the transformer-based model for 3D instance segmentation proposed in [Schult23ICRA], designed to iteratively attend to hierarchical feature representations. The model processes the 4 coarsest levels of a ResNet-based U-Net backbone across three iterations, progressively refining from coarse to fine, resulting in $L=12$ transformer decoder layers. Each transformer decoder layer shares weights across iterations and consists of a standard transformer layer utilizing self-attention and masked cross-attention mechanisms. The feature backbone employed is Minkowski Res16UNet34C [choy20194d].
Training Details. We adopt the data augmentation, hyperparameters, and training strategy described in [Schult23ICRA]. For joint training (Row 1, Table 4), the model is trained for 600 epochs using the AdamW optimizer [loshchilov2017decoupled] with a one-cycle learning rate scheduler [smith2018superconvergence], and results are evaluated on the entire validation set. In incremental training, we retain the same hyperparameters, adjusting only the number of epochs, and we use a memory buffer size of 50 scenes. Training is divided into three phases, introducing one split per phase across the three designed scenarios. After each phase, the resulting model is evaluated on all classes encountered up to that point.
5.2 Results and Discussion
Table 1: Comparison between the baseline and proposed method with mAP25, mAP50, and mAP, which is after training for all the stages. We also report the FPP metric.
Scenarios Methods Average Precision $\uparrow$ FPP $\downarrow$
mAP25 mAP50 mAP mAP25 mAP50
Baseline 16.46 14.29 10.44 51.30 46.82
Split_A CLIMB-3D 35.69 31.05 22.72 3.44 2.63
Baseline 17.22 15.07 10.93 46.27 42.1
Split_B CLIMB-3D 35.48 31.56 23.69 8.00 5.51
Baseline 25.65 21.08 14.85 31.68 28.84
Split_C CLIMB-3D 31.59 26.78 18.93 9.10 7.89
To evaluate our proposed method, we conduct a comparative analysis using exemplar replay (ER) for instance segmentation and [Yang_2023_CVPR] for semantic segmentation as baselines. As shown in Table 1, our method, which integrates exemplar replay, knowledge distillation, and an imbalance correction (IC) module, achieves notable improvements over the baseline in terms of mAP and FPP. Specifically, in the Split_A  scenario, our approach significantly enhances overall performance. We observe an improvement of 19.23%, 31.05%, and 12.28% for $\text{mAP}_{50}$, $\text{mAP}_{25}$, and overall mAP, respectively, while reducing forgetting by 47.86% as measured by $\text{mAP}_{50}$. For the Split_B  scenario as well, CLIMB-3D demonstrates a consistent performance boost over the baseline, significantly improving mAP and reducing forgetting, which is lowered to 5.52% compared to 46.21% in the baseline for AP50. Likewise, in the Split_C  scenario, CLIMB-3D enhances both learning efficiency and forgetting reduction, achieving a performance of 26.78% in $\text{mAP}_{50}$  and reducing forgetting by 20.95% compared to baseline. These results across scenarios underscore the effectiveness of our approach.
Table 2: Comparison with previous method on semantic segmentation on ScanNet V2 dataset. We report the mIoU metric for the evaluated models.
Methods Phase=1 Phase=2 All
EWC [serra2018overcoming] 17.75 13.22 16.62
LwF [li2017learning] 30.38 13.37 26.13
Yang et al. [Yang_2023_CVPR] 34.16 13.43 28.98
CLIMB-3D (Ours) 69.39 32.56 59.38
Although our method focuses on segmenting individual objects (instance segmentation), we also demonstrate its performance in semantic segmentation by presenting a comparative analysis with existing methods for class-incremental semantic segmentation on the ScanNet V2 dataset (Table 2). Using our predicted labels, we assign each point the label corresponding to the highest confidence mask and exclude background labels (floor and wall), as these are not part of the object-level segmentation. Following the dataset splits established by [Yang_2023_CVPR], we report results for both training phases. Our proposed method achieves a substantial improvement over prior methods, with a gain of 35.23% in Phase 1 and approximately 19.1% in Phase 2. Overall, our method reaches a mIoU of 59.38%, significantly outperforming previous baselines, which achieve a lower mIoU of around 30%.
Table 3: Comparison of results in terms of mAP50 with and proposed CLIMB-3D for three different scenarios. Each scenario is trained in three phases ($\mathtt{phase}=1,2,3$) by introducing a single split s at a time. The results highlighted in orange are with the proposed method, and the best results for each scenario are in bold.
Scenarios Methods phase=1 phase=2 phase=3
s1 s1 s2 All s1 s2 s3 All
Baseline 56.82 18.51 32.81 25.72 10.38 9.43 24.27 14.28
Split_A CLIMB-3D 56.82 54.67 33.75 44.13 54.19 12.02 26.55 31.05
Baseline 51.57 13.32 42.21 24.53 9.55 12.45 26.78 15.07
Split_B CLIMB-3D 51.57 46.74 37.45 43.13 46.06 15.95 26.68 31.56
Baseline 36.40 7.74 37.62 22.32 7.55 15.96 40.41 21.08
Split_C CLIMB-3D 36.40 32.63 33.38 33.00 28.51 17.11 34.64 26.78
We extend the analysis from Table 1 to Table 3 to highlight the impact of our proposed method on individual splits across various scenarios. The results clearly demonstrate that our model consistently retains knowledge of previous tasks better than the baseline. For Split_A , our model shows improvement throughout the phase. In Phase 3 of (s2), although both the baseline and our method exhibit a performance drop, our method reduces forgetting significantly compared to the baseline. The Split_B  scenario, while more complex than Split_A, achieves comparable results due to semantic similarity among classes within the same task. In Phase 2, our model achieves overall all 43.13% $\text{mAP}_{50}$ compared to 24.53% on baseline, a similar trend is observed in Phase 3, where our method not only consistently improves learning but also enhances retention of previous information. After all three tasks, our method achieves an overall performance of 31.56% AP50, compared to 15.07% for the baseline. In the Split_C  scenario, the first-stage model struggles due to the increased complexity introduced by random grouping. In Phase 2, while the baseline focuses on learning the current task, it suffers from severe forgetting of prior knowledge. Conversely, our method balances new task learning with the retention of earlier information. By Phase 3, the model effectively consolidates s1 and maintains strong performance across all task splits. Overall, our proposed method improves mAP by 5.6%.
5.3 Ablation
To assess the effectiveness of each component in our proposed framework, we perform an ablation study. Initially, we establish an upper-bound performance by jointly training the model on the complete dataset using a transformer-based architecture, such as Mask3D [Schult23ICRA], referred to as the Oracle. For the incremental learning setup, we generate training splits according to the scenarios outlined earlier. In this study, we first train the model naively across phases and then sequentially integrate each module to evaluate its individual contribution to performance. Table 4 summarizes the results for the Split_A  scenario, using both the $\text{mAP}_{50}$  and FPP metrics.
Table 4: Ablation study results illustrating the impact of exemplar replay, knowledge distillation, and imbalance correction modules in a three-phase training setup. Each split, representing the subset of data introduced at each phase (p), is labeled as ‘s’ followed by the phase number. The final column, ‘All’, in each phase reports performance across all classes encountered up to that phase. Joint training results (Oracle) are highlighted in gray, while results with all modules combined are marked in orange. The best-performing results are shown in bold.
Row Modules p=1 $\uparrow$ p=2 $\uparrow$ p=3 $\uparrow$ FPP $\downarrow$
s1 s1 s2 All s1 s2 s3 All
1. Oracle - - - - 55.14 30.77 25.30 37.68 -
2. Naïve 56.82 0.00 28.09 14.15 0.00 0.00 19.67 5.80 56.82
3. + ER 56.82 18.51 32.81 25.72 10.38 9.43 24.27 14.28 46.44
4. + KD 56.82 50.00 34.39 42.13 49.78 11.41 26.47 29.28 7.04
5. + IC 56.82 54.67 33.75 44.13 54.19 12.02 26.55 31.05 2.63
Naïve Training. In the naive incremental training setup, where no dedicated modules are incorporated, the model learns the current task but suffers from catastrophic forgetting of the previously learned tasks, as expected. This behavior is evident in row 2, where, upon transitioning to phase 2, the model entirely forgets the classes learned during phase 1. A similar trend is observed in phase 3, and this pattern is also reflected in the FPP metric.
Effect of Exemplar Replay. To mitigate catastrophic forgetting, we incorporate exemplar replay, which stores and replays examples from previous tasks. As shown in row 3, exemplar replay improves average precision by 18.5% for s1 in phase 2 and 10.38% in phase 3. It also reduces forgetting for s2 by 9.43% in phase 3, while slightly improving learning on the current task. However, substantial forgetting persists, as reflected in the FPP metric, highlighting the limitations of exemplar replay alone.
Effect of Knowledge Distillation. The addition of knowledge distillation (KD), which retains a copy of the model from previous tasks, facilitates the preservation of past task knowledge while enabling forward knowledge transfer. As shown in row 4, KD considerably reduces forgetting and boosts performance on the current task. Specifically, for s1, KD improves $\text{mAP}_{50}$ by 31.49% in phase 2 and by 39.40% in phase 3, compared to exemplar replay. Overall, KD leads to a 15.00% increase in performance while reducing forgetting by 39.40% after all tasks have been learned.
Effect of Imbalance Correction. The imbalance correction module addresses the class imbalance in the dataset by re-weighting the teacher model’s predictions during KD based on class frequency. As highlighted in row 5 of Table 4, this addition further improves performance. Specifically, for s1, imbalance correction reduces forgetting by 4.67% and 4.41% in phases 2 and 3, respectively, compared to the results without this module (row 4). For s2, while a slight decrease in current task performance is observed in phase 2, this is likely due to the module’s prioritization of mitigating forgetting less frequent classes in previous tasks. In phase 3, performance on s2 and s3 improves. Overall, imbalance correction significantly reduces forgetting, achieving improvements of 4.41% and 43.81% over KD and exemplar replay, respectively.
6 Conclusion
We address the challenge of class-incremental 3D instance segmentation with class imbalance. We propose an innovative approach that integrates a memory-efficient exemplar replay buffer, knowledge distillation, and a novel imbalance correction module. This framework mitigates the forgetting of rare classes during incremental learning by accounting for the frequency of object occurrences. To enable comprehensive evaluation, we design three incremental learning scenarios, each comprising three phases that reflect real-world dynamics. Our experimental results demonstrate that the proposed framework significantly enhances the learning of new classes while reducing forgetting of previously learned ones. The carefully designed scenarios and framework not only offer a strong baseline but also provide a clear benchmark for future research, laying a foundation for more advanced techniques in class-incremental learning.
Supplementary Material
In this supplementary material, we first demonstrate the performance gains on rare classes achieved by incorporating the IC module in Appendix A. Next, we provide detailed split information for all scenarios, based on class names, in Appendix B. Finally, we present a qualitative comparison between the baseline method and our proposed approach in Appendix C.
Appendix Appendix A Evaluation on Rare Categories
The proposed imbalance correction (IC) module, as detailed in Section 4.2, is designed to address the performance gap for rare classes. To assess its impact, we compare its performance with the framework which has exemplar replay (ER) and knowledge distillation (KD). Specifically, we focus on its ability to improve performance for rare classes, which the model encounters infrequently compared to more common classes.
Table 5: Results for classes observed by the model 1–20 times during an epoch, evaluated on Split_A  for Phase 2, in terms of $\text{mAP}_{50}$.
Classes Seen Count ER+KD ER+KD+IC
paper towel dispenser 2 73.10 74.90
recycling bin 3 55.80 60.50
ladder 5 53.90 57.10
trash bin 7 31.50 57.30
bulletin board 8 23.30 38.20
shelf 11 48.00 50.50
dresser 12 44.00 55.80
copier 12 93.30 94.50
object 12 3.10 3.30
stairs 13 51.70 67.70
bathtub 16 80.30 86.60
oven 16 1.50 3.30
divider 18 36.40 45.00
column 20 57.30 75.00
Average - 46.66 54.98
The results, shown in Table 5 and Table 6, correspond to evaluations on Split_A  for Phase 2 and Phase 3, respectively. In Phase 2, we evaluate classes seen 1–20 times per epoch, while Phase 3 targets even less frequent classes, with observations limited to 1–10 times per epoch.
As illustrated in Table 5, the IC module substantially improves performance on rare classes in terms of $\text{mAP}_{50}$  in Phase 2 of Split_A. For instance, classes like recycling bin and trash bin, seen only 3 and 7 times, respectively, shows significant improvement when the IC module is applied. Overall, the IC module provides an average boost of 8.32%, highlighting its effectiveness in mitigating class imbalance.
Table 6: Results for classes observed by the model 1–10 times during an epoch, evaluated on Split_A  for Phase 3, in terms of $\text{mAP}_{50}$.
Classes Seen Count ER+KD ER+KD+IC
piano 1 7.10 59.40
bucket 1 21.10 31.50
laundry basket 1 3.80 17.40
dresser 2 55.00 55.40
paper towel dispenser 2 32.50 35.50
cup 2 24.70 30.30
bar 2 35.40 39.50
divider 2 28.60 42.40
case of water bottles 2 0.00 1.70
shower 3 0.00 45.50
mirror 8 56.00 68.80
trash bin 4 1.10 2.70
backpack 5 74.50 76.70
copier 5 94.00 96.80
bathroom counter 3 3.90 20.30
ottoman 4 32.60 36.20
storage bin 3 5.10 10.50
dishwasher 3 47.40 66.20
trash bin 4 1.10 2.70
backpack 5 74.50 76.70
copier 5 94.00 96.80
sofa chair 6 14.10 43.50
file cabinet 6 49.20 57.60
tv stand 7 67.70 68.60
mirror 8 56.00 68.80
blackboard 8 57.10 82.80
clothes dryer 9 1.70 3.20
toaster 9 0.10 25.90
wardrobe 10 22.80 58.80
jacket 10 1.20 4.10
Average - 32.08 44.21
Similarly, Table 6 presents results for Phase 3, demonstrating significant gains for infrequent classes. For example, even though the classes such as piano, bucket, and laundry basket are observed only once, IC module improves the performance by 52.30%, 10.40%, and 13.60%, respectively. The ER+KD module does not focus on rare classes like shower and toaster which results in low performance, but the IC module compensates for this imbalance by focusing on underrepresented categories. On average, the addition of the proposed IC module into the framework outperforms ER+KD by 12.13%.
Table 7: Classes grouped by tasks for each proposed scenario on the ScanNet200 dataset labels. The three scenarios Split_A, Split_A, and Split_C  are each divided into three tasks: Task 1, Task 2, and Task 3.
Split_A Split_B Split_C
Task 1 Task 2 Task 3 Task 1 Task 2 Task 3 Task 1 Task 2 Task 3
chair wall pillow tv stand cushion paper broom fan rack
table floor picture curtain end table plate towel stove music stand
couch door book blinds dining table soap dispenser fireplace tv bed
desk cabinet box shower curtain keyboard bucket blanket dustpan soap dish
office chair shelf lamp bookshelf bag clock dining table sink closet door
bed window towel tv toilet paper guitar shelf toaster basket
sink bookshelf clothes kitchen cabinet printer toilet paper holder rail doorframe chair
toilet curtain cushion pillow blanket speaker bathroom counter wall toilet paper
monitor kitchen cabinet plant lamp microwave cup plunger mattress ball
armchair counter bag dresser shoe paper towel roll bin stand monitor
coffee table ceiling backpack monitor computer tower bar armchair copier bathroom cabinet
refrigerator whiteboard toilet paper object bottle toaster trash bin ironing board shoe
tv shower curtain blanket ceiling bin ironing board dishwasher radiator blackboard
nightstand closet shoe board ottoman soap dish lamp keyboard vent
dresser computer tower bottle stove bench toilet paper dispenser projector toaster oven bag
stool board basket closet wall basket fire extinguisher potted plant paper bag paper
bathtub mirror fan couch fan ball coat rack structure projector screen
end table shower paper office chair laptop hat end table picture pillar
dining table blinds person kitchen counter person shower curtain rod tissue box purse range hood
keyboard rack plate shower paper towel dispenser paper cutter stairs tray coffee maker
printer blackboard container closet oven tray fire extinguisher couch handicap bar
tv stand rail soap dispenser doorframe rack toaster oven case of water bottles telephone pillow
trash can radiator telephone sofa chair piano mouse water bottle shower curtain rod decoration
stairs wardrobe bucket mailbox suitcase toilet seat cover dispenser ledge trash can printer
microwave column clock nightstand rail storage container shower head closet wall object
stove ladder stand washing machine container scale guitar case cart mirror
bin bathroom stall light picture telephone tissue box kitchen cabinet hat ottoman
ottoman shower wall pipe book stand light switch poster paper cutter water pitcher
bench mat guitar sink light crate candle storage organizer refrigerator
washing machine windowsill toilet paper holder recycling bin laundry basket power outlet bowl vacuum cleaner divider
copier bulletin board speaker table pipe sign plate mouse toilet
sofa chair doorframe bicycle backpack seat projector person paper towel roll washing machine
file cabinet shower curtain rod cup shower wall column candle storage bin laundry detergent mat
laptop paper cutter jacket toilet bicycle plunger microwave calendar scale
paper towel dispenser shower door paper towel roll copier ladder stuffed animal office chair wardrobe dresser
oven pillar machine counter jacket headphones clothes dryer whiteboard bookshelf
piano ledge soap dish stool storage bin broom headphones laundry basket tv stand
suitcase light switch fire extinguisher refrigerator coffee maker guitar case toilet seat cover dispenser shower door closet rod
recycling bin closet door ball window dishwasher dustpan bathroom stall door curtain plant
laundry basket shower floor hat file cabinet machine hair dryer speaker folded chair counter
clothes dryer projector screen water cooler chair mat water bottle keyboard piano suitcase bench
seat divider mouse wall windowsill handicap bar cushion hair dryer ceiling
storage bin closet wall scale plant bulletin board purse table mini fridge piano
coffee maker bathroom stall door power outlet coffee table fireplace vent nightstand dumbbell closet
dishwasher stair rail decoration stairs mini fridge shower floor bathroom vanity oven cabinet
bar bathroom cabinet sign armchair water cooler water pitcher laptop luggage cup
toaster closet rod projector cabinet shower door bowl shower wall bar laundry hamper
ironing board structure vacuum cleaner bathroom vanity pillar paper bag desk pipe light switch
fireplace coat rack candle bathroom stall ledge alarm clock computer tower bathroom stall cd case
kitchen counter storage organizer plunger mirror furniture music stand soap dispenser blinds backpack
toilet paper dispenser stuffed animal blackboard cart laundry detergent container toilet paper dispenser windowsill
mini fridge headphones trash can decoration dumbbell bicycle coffee table box
tray broom stair rail closet door tube light dish rack book
toaster oven guitar case box vacuum cleaner cd case clothes guitar mailbox
toilet seat cover dispenser hair dryer towel dish rack closet rod machine seat sofa chair
furniture water bottle door range hood coffee kettle furniture clock shower curtain
cart purse clothes projector screen shower head stair rail alarm clock bulletin board
storage container vent whiteboard divider keyboard piano toilet paper holder board crate
tissue box water pitcher bed bathroom counter case of water bottles floor file cabinet tube
crate bowl floor laundry hamper coat rack bucket ceiling light window
dish rack paper bag bathtub bathroom stall door folded chair stool ladder power outlet
range hood alarm clock desk ceiling light fire alarm door paper towel dispenser power strip
dustpan laundry detergent wardrobe trash bin power strip sign shower floor bathtub
handicap bar object clothes dryer bathroom cabinet calendar recycling bin stuffed animal column
mailbox ceiling light radiator structure poster shower water cooler fire alarm
music stand dumbbell shelf storage organizer luggage jacket coffee kettle storage container
bathroom counter tube potted plant bottle kitchen counter
bathroom vanity cd case mattress
laundry hamper coffee kettle
trash bin shower head
keyboard piano case of water bottles
folded chair fire alarm
luggage power strip
mattress calendar
poster
potted plant
Appendix Appendix B Incremental Scenarios Phases
Table 7 presents the task splits for each proposed scenario introduced in Section 4.3 using the ScanNet200 dataset. The three scenarios, Split_A, Split_B, and Split_C, are each divided into three tasks: Task 1, Task 2, and Task 3. Notably, the order of classes in these tasks is random.
Appendix Appendix C Qualitative Results
In this section, we present a qualitative comparison of the proposed framework with the baseline method. Figure 4 illustrates the results on the Split_A evaluation after learning all tasks, comparing the performance of the baseline method and our proposed approach. As shown in the figure, our method demonstrates superior instance segmentation performance compared to the baseline. For example, in row 1, the baseline method fails to segment the sink, while in row 3, the sofa instance is missed. Overall, our framework consistently outperforms the baseline, with several missed instances by the baseline highlighted in red circles.
Figure 4: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split_A  evaluation after learning all tasks.
In Figure 5, we present the results on Split_B, highlighting instances where the baseline method underperforms, marked with red circles. For example, in row 2, the baseline method incorrectly identifies the same sofa as separate instances. Similarly, in row 5, the washing machine is segmented into two instances by the baseline. In contrast, the proposed method delivers results that closely align with the ground truth, demonstrating its superior performance
Figure 5: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split_B  evaluation after learning all tasks.
Similarly, Figure 6 highlights the results on Split_C, where classes are encountered in random order. The comparison emphasizes the advantages of our method, as highlighted by red circles. The baseline method often misses instances or splits a single instance into multiple parts. In contrast, our approach consistently produces results that are closely aligned with the ground truth, further underscoring its effectiveness.
Figure 6: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split_C  evaluation after learning all tasks.

### v2

@@
 Abstract
-While 3D instance segmentation has made significant progress, current methods struggle to address realistic scenarios where new categories emerge over time with natural class imbalance.
-This limitation stems from existing datasets, which typically feature few well-balanced classes.
-Although few datasets include unbalanced class annotations, they lack the diverse incremental scenarios necessary for evaluating methods under incremental settings.
-Addressing these challenges requires frameworks that handle both incremental learning and class imbalance.
-However, existing methods for 3D incremental segmentation rely heavily on large exemplar replay, focusing only on incremental learning while neglecting class imbalance.
-Moreover, frequency-based tuning for balanced learning is impractical in these setups due to the lack of prior class statistics.
-To overcome these limitations, we propose a framework to tackle both Continual Learning and class Imbalance for 3D instance segmentation (CLIMB-3D).
-Our proposed approach combines Exemplar Replay (ER), Knowledge Distillation (KD), and a novel Imbalance Correction (IC) module.
-Unlike prior methods, our framework minimizes ER usage, with KD preventing forgetting and supporting the IC module in compiling past class statistics to balance learning of rare classes during incremental updates.
-To evaluate our framework, we design three incremental scenarios based on class frequency, semantic similarity, and random grouping that aim to mirror real-world dynamics in 3D environments.
-Experimental results show that our proposed framework achieves state-of-the-art performance, with an increase of up to 16.76% in mAP compared to the baseline.
-Code is available at: https://github.com/vgthengane/CLIMB3D
+While 3D instance segmentation (3DIS) has advanced significantly, existing methods typically assume that all object classes are known in advance and are uniformly distributed.
+However, this assumption is unrealistic in dynamic, real-world environments where new classes emerge gradually and exhibit natural imbalance.
+Although some approaches have addressed class emergence, they often overlook class imbalance, resulting in suboptimal performance – particularly on rare categories.
+To tackle this challenge, we propose CLIMB-3D, a unified framework for CLass-incremental Imbalance-aware 3DIS.
+Building upon established exemplar replay (ER) strategies, we show that ER alone is insufficient to achieve robust performance under constrained memory conditions.
+To mitigate this, we introduce a novel pseudo-label generator (PLG) that extends supervision to previously learned categories by leveraging predictions from a frozen prior model.
+Despite its promise, PLG tends to bias towards frequent classes.
+Therefore, we propose a class-balanced re-weighting (CBR) scheme, that estimates object frequencies from pseudo-labels and dynamically adjusts training bias – without requiring access to past data.
+We design and evaluate three incremental scenarios for 3DIS on the challenging ScanNet200 dataset, and additionally on semantic segmentation on ScanNetV2.
+Our approach achieves state-of-the-art results, surpassing prior work by up to 16.76% mAP for instance segmentation and approximately 30% mIoU for semantic segmentation, demonstrating strong generalization across both frequent and rare classes.
 1 Introduction
-Figure 1: Class-incremental setup for 3D instance segmentation.
-As tasks progress through time, new classes are introduced incrementally.
-After each new task, the model should recognize both previously learned and newly introduced classes.
-For example, at Task-2, new classes such as Pillow, Coffee Table, and Sofa Chair are added, and the model is able to detect these three classes along with previous ones like Table, Chair, and Couch.
-3D instance segmentation is an essential task in computer vision that involves identifying and segmenting individual objects in the real physical space, playing a key role in applications across graphics, robotics, and autonomous systems.
-Its ability to provide precise object boundaries and class labels enhances scene understanding, facilitates object manipulation, and improves perception in dynamic environments.
-In recent years, a variety of methods have been proposed, including top-down approaches [wang2018sgpn, jia2021scaling, zhang2021point], bottom-up approaches [hou20193dsis, yang2019learning], and transformer-based architectures [Schult23ICRA].
-These methods have shown impressive results in traditional setups, which assume that all object classes are available during training.
-However, this assumption limits its applicability in real-world scenarios where new categories gradually emerge over time, often exhibiting naturally imbalanced distributions.
-Thus, there is a need for class-incremental learning (CIL) frameworks that not only adapt to new classes but also preserve prior knowledge, especially for rare or less frequent categories, which are more prone to catastrophic forgetting.
-Most existing research in class-incremental learning focuses on 2D image classification [rebuffi2017icarl, li2017learning, aljundi2018memory, serra2018overcoming], with some extensions to object detection [liu2023continual, joseph2021towards, shmelkov2017incremental] and semantic segmentation [cermelli2020modeling, cermelli2022incremental, douillard2021plop].
-These methods employ strategies such as exemplar replay [buzzega2020dark, rebuffi2017icarl, cha2021co2l, kamra2017deep], regularization [aljundi2018memory, li2017learning, serra2018overcoming], and knowledge distillation [douillard2020podnet, kang2022class] to preserve previously learned knowledge and mitigate catastrophic forgetting [mccloskey1989catastrophic].
-Few studies have applied CIL to point clouds; however, they mostly focus on object-level classification [dong2021i3dol, liu2021l3doc, chowdhury2021learning].
-At the scene level, some works have explored 3D semantic segmentation [Yang_2023_CVPR] with incremental learning, but their performance is not as competitive as state-of-the-art methods that do not employ incremental learning, which limits their applicability.
-Other methods tackle open-world incremental learning [boudjoghra20243d] but rely heavily on large exemplar sets [rolnick2019experience] and often neglect the challenge of class imbalance.
-To address this, we propose CLIMB-3D, a unified framework that combines Exemplar Replay (ER), Knowledge Distillation (KD), and a novel Imbalance Correction (IC) module to tackle Continual Learning for Imbalance 3D instance segmentation in indoor environments, as shown in Figure 1.
-Our framework operates as follows: ER stores a subset of representative samples from previous stages, allowing for effective replay during new task learning.
-KD transfers knowledge by retaining a copy of the model from the previous task, thereby mitigating forgetting.
-The IC module is specifically designed to reduce forgetting of rare classes by leveraging the frequency of object occurrence.
-However, since we do not have access to the previous task data or statistics during incremental phases, we instead compile these statistics from the previous model used by KD to generate weight for earlier categories.
-These weights are used to favor both frequent and rare classes, ensuring a balanced learning and mitigating forgetting.
-To evaluate CLIMB-3D in a realistic, incremental learning setup, we create three benchmark scenarios using the ScanNet200 dataset [rozenberszki2022language], which features 200 classes with natural class imbalances.
-These scenarios are designed to reflect real-world conditions where new categories emerge gradually and follow inherent class imbalances.
-These are based on 1 frequency of object occurrence, 2 semantic similarity between the object, and 3 random grouping.
-Our experiments demonstrate that CLIMB-3D significantly improves performance by effectively mitigating the forgetting of previous tasks compared to earlier exploration in class incremental 3D segmentation.
-In summary, our contributions are as follows:
-  1. 1.
-    We propose a new problem setting for imbalanced class incremental 3D segmentation, along with a simple yet effective method to address this challenge by minimizing catastrophic forgetting and balancing the learning process.
-  2. 2.
-    To benchmark this setting, we design three scenarios aimed at simulating real-world conditions where objects emerge continuously with natural class imbalance.
-  3. 3.
-    Experimental results show that our proposed framework achieves state-of-the-art performance, with an increase of up to 16.76% in mAP compared to the baseline.
+Figure 1: Overview of the CI-3DIS setting: new classes are introduced incrementally as tasks progress.
+After each task, the model must recognize both previously learned and newly added classes.
+For instance, at Task 2, classes like Pillow, Coffee Table, and Sofa Chair are introduced, while the model retains the ability to detect earlier classes such as Table, Chair, and Couch.
+3D instance segmentation (3DIS) is an essential task in computer vision with widespread applications in graphics, robotics, and augmented reality [boudjoghra20243d, lai2022stratified].
+It involves identifying and segmenting individual objects in 3D space, enabling precise object boundaries and class labels that enhance scene understanding and perception in dynamic environments.
+In recent years, a variety of methods for 3DIS have been proposed, including top-down [wang2018sgpn, jia2021scaling, zhang2021point], bottom-up [hou20193dsis, yang2019learning], and transformer-based approaches [Schult23ICRA].
+These methods have demonstrated strong performance under the traditional setting, which assumes that all object classes are available and well-balanced during training.
+However, these assumptions limit their applicability in real world settings, where new categories gradually emerge over time, often exhibiting natural class imbalance.
+These limitations in current 3DIS methods underscore the need for a Class-Incremental Learning (CIL) [rebuffi2017icarl, de2021continual] framework capable of learning new categories as they appear, while maintaining accuracy on previously learned classes – especially rare ones, which are more susceptible to without catastrophic forgetting [mccloskey1989catastrophic].
+Unfortunately, most existing research in CIL has focused on the 2D domain for image classification [rebuffi2017icarl, li2017learning, aljundi2018memory, serra2018overcoming], with limited extensions to object detection [liu2023continual, joseph2021towards, shmelkov2017incremental] and semantic segmentation [cermelli2020modeling, cermelli2022incremental, douillard2021plop].
+Extensions to 3D point clouds remain limited, with most work focusing on object-level classification rather than scene-level segmentation [dong2021i3dol, liu2021l3doc, chowdhury2021learning].
+Recent works on scene-level class-incremental 3DIS (CI-3DIS) [boudjoghra20243d] and class-incremental semantic segmentation (CI-3DSS) [Yang_2023_CVPR] show promise, but typically rely on large exemplar buffers [rolnick2019experience] Moreover, they typically overlook class imbalance, limiting real-world applicability.
+To address this, we propose CLIMB-3D, a unified framework for CLass-incremental IMBalance-aware 3DIS that jointly tackles catastrophic forgetting and class imbalance.
+The framework begins with Exemplar Replay (ER), storing a small number of samples from past classes for later replay.
+However, this alone does not yield promising results under the strict memory constraints needed CI-3DIS.
+To address this, we introduce a Pseudo-Label Generator (PLG), which uses a frozen previous task model to generate supervision for earlier classes.
+However, we observed that, PLG tends to favour frequent classes while ignoring under-represented ones.
+To mitigate this, we propose a Class-Balanced Re-weighting (CBR), which estimates class frequencies from the pseudo-labels to enhance the learning of rare classes without accessing past data.
+Together, these components form a practical and effective solution for real-world 3DIS.
+To evaluate CLIMB-3D in a realistic incremental setup, we design three benchmark scenarios using the ScanNet200 [rozenberszki2022language] dataset, which contains 200 classes with a natural long-tail distribution.
+These scenarios simulate real-world conditions where new classes gradually emerge, following inherently imbalanced distributions.
+Specifically, new classes appear in groups based on one of three criteria: (A) object frequency, (B) semantic similarity, or (C) random grouping without a specific pattern.
+Our experiments on these scenarios for incremental 3DIS show that CLIMB-3D effectively mitigates forgetting of previous tasks while improving performance across both frequent and rare classes, delivering consistent gains over prior methods.
+In summary, our contributions are threefold: (i) We introduce a novel problem setting of imbalanced class-incremental 3DIS, along with a simple and effective method to mitigate catastrophic forgetting and ensure balanced learning; (ii) We establish three benchmark scenarios that realistically model continuous object emergence with natural class imbalance; and (iii) We validate our framework’s effectiveness, achieving SOTA results with improvements up to 16.76% mAP over existing baselines [^1] Code will be available at: https://github.com/vgthengane/CLIMB3D .
 2 Related Work
@@
 In contrast, our work introduces a dedicated continual learning framework for 3D instance segmentation, with a focus on effective knowledge transfer from previous tasks, while also accounting for the challenges posed by infrequent class occurrences.
-3 Preliminaries
-3.1 Transformer-based Segmentation
-We adopt a transformer-based instance segmentation method based on Mask3D [Schult23ICRA].
-Specifically, transformer-based segmentation model $\Phi$ is employed for point cloud instance segmentation.
-Given an input point cloud $p$, the model predicts $\hat{y}=\{(\hat{m_{j}},\hat{c_{j}})\}_{j=1}^{J}$, which consists of mask predictions and class probabilities for each instance.
-The segmentation process begins by quantizing the input point cloud $p$ into voxels $V$, creating voxelized representations of size $\mathbb{R}^{M_{0}\times 3}$.
-Each voxel is assigned an average RGB color computed from the points within that voxel, serving as its initial feature representation.
-The feature backbone network generates a high-resolution output feature map $\mathbf{F}_{0}\in\mathbb{R}^{M_{0}}$.
-Additionally, intermediate feature maps are extracted from the decoder layers of the backbone network.
-For each intermediate feature map ($r\geq 0$), a set of $K_{r}$ voxels is selected, and their features are linearly projected to a fixed dimension $D$, yielding feature maps $\mathbf{F}_{r}\in\mathbb{R}^{M_{r}\times D}$.
-The Transformer decoder initiates with a set of K instance queries and iteratively improves them using L Transformer decoder layers.
-These layers employ cross-attention to refine the instance queries, incorporating information from point cloud features.
-The decoder attends to a specific feature map obtained from the corresponding feature backbone layer at each layer, employing conventional cross-attention mechanisms.
-This process enables the decoder to reason at the instance level through self-attention, resulting in the generation of accurate and contextually relevant instance queries tailored to the specific scene.
-To achieve this, the voxel features $\mathbf{F}_{r}\in\mathbb{R}^{M_{r}\times D}$ are transformed into sets of keys $\mathbf{K}\in\mathbb{R}^{M_{r}\times D}$ and values $\mathbf{V}\in\mathbb{R}^{M_{r}\times D}$ through linear projection.
-The instance queries $\mathbf{Z}$ are also projected to create the queries $\mathbf{Q}$.
-This enables cross-attention, allowing the queries to gather relevant information from the voxel features.
-Following cross-attention, a self-attention step occurs among the queries, facilitating information exchange and refinement.
-The learned queries are then used to make $K$ class and mask predictions, which are matched with ground truth labels through bipartite matching, resulting in $\hat{y}=\{(\hat{m}_{j},\hat{c}_{j})\}_{j=1}^{J}$.
-The model is optimized based on the ground truth label, mask, and class predictions:
-$\mathcal{L}_{\mathrm{Seg}}(y_{j},\hat{y}_{j})=\mathcal{L}_{\mathrm{mask}}(m_{j},\hat{m}_{j})+\lambda_{\mathrm{cls}}\mathcal{L}_{\mathrm{cls}}(c_{j},\hat{c}_{j})$ (1)
-where, mask loss $\mathcal{L}_{\mathrm{mask}}=\lambda_{\mathrm{ce}}\mathcal{L}_{\mathrm{ce}}(y_{j},\hat{y}_{j})+\lambda_{\mathrm{dice}}\mathcal{L}_{\mathrm{dice}}(y_{j},\hat{y}_{j})$ and $\lambda_{\mathrm{cls}}\mathcal{L}_{\mathrm{cls}}$ is classification loss.
-The traditional setup assumes all categories are available and well-balanced during training.
-However, in scenarios where only a subset of categories is present, training the model in multiple phases is required.
-Unfortunately, such multi-phase models often suffer from forgetting previous tasks.
-To address this issue, we employ incremental learning strategies which will be discussed in the next section.
 3 Methodology
-In this section, we begin by formulating the incremental 3D instance segmentation problem and introduce our proposed method to address it using a transformer-based model.
-We then detail the design of incremental scenarios, outlining the motivation and considerations behind each one.
-Figure 2: CLIMB-3D: At phase $t$ (for $t>1$), we create a copy of the model from the previous phase, $\Phi^{\mathrm{T}}$ (the teacher model), and designate it as $\Phi^{\mathrm{S}}$ (the student model).
-Both $\Phi^{\mathrm{T}}$ and $\Phi^{\mathrm{S}}$ process the input point cloud simultaneously, producing predictions $\hat{y}_{\mathrm{pseudo}}$ and $\hat{y}$, respectively.
-To prevent the model from forgetting less frequent categories from previous tasks, we balance the predictions from $\Phi^{\mathrm{T}}$, resulting in a balanced output $\hat{y}^{\mathrm{(w)}}_{\mathrm{pseudo}}$.
-The concatenated vector of ground truth labels $[y,\hat{y}_{\mathrm{pseudo}},\hat{y}^{\mathrm{(w)}}_{\mathrm{pseudo}}]$ is then compared with the predicted labels $\hat{y}$.
-A loss function is applied based on this comparison, enabling the student model to learn from the differences between the predicted and the concatenated pseudo-ground truth labels
+Figure 2: Overview of CLIMB-3D for CI-3DIS.
+The model incrementally learns new classes in sequential phases using ER to mitigate forgetting, a PLG for supervision on old classes, and CBR to handle class imbalance.
 3.1 Problem Formulation
-The objective of 3D point cloud instance segmentation is to accurately identify and segment individual instances of objects within a given point cloud.
-Mathematically, the training dataset is represented as $\mathcal{D}=(\mathcal{P},\mathcal{Y})=\{(p_{i},y_{i})\}_{i=1}^{N}$, where $N$ is the total number of samples.
-Each sample consists of a colored point cloud $p_{i}\in\mathbb{R}^{M\times 6}$ of size $M$, where the point coordinates and color values are represented as $\{x,y,z,r,g,b\}$.
-The corresponding annotations are denoted as $y_{i}=\{(m_{i,j},c_{i,j})\}_{j=1}^{J}$, where $m_{i,j}$ represents the instance mask for the $j$-th instance, and $c_{i,j}\in\mathcal{C}=\{1,\dots,C\}$ denotes the semantic label of the object category to which the instance belongs for the $i$-th point cloud.
-Here, $J$ represents the total number of instances in the $i$-th point cloud, and $C$ indicates the number of distinct object categories.
-During the learning process, the model $\Phi$ will process this dataset and output predictions $\hat{y}_{i,j}=(\hat{m}_{i,j},\hat{c}_{i,j})$, where $\hat{m}_{i,j}$ represents the predicted instance mask and $\hat{c}_{i,j}$ denotes the predicted semantic label for the $j$-th instance in the $i$-th point cloud.
-To adapt the dataset to an incremental learning setting, we partition the object categories $\mathcal{C}$ into $T$ subsets, denoted as $\mathcal{C}=\mathcal{C}^{1}\cup\cdots\cup\mathcal{C}^{T}$.
-Each phase $t\in\{1,\dots,T\}$ is associated with a specific subset $\mathcal{C}^{t}$, and its corresponding dataset $\mathcal{D}^{t}$ which only contains annotations for objects belonging to the corresponding subset.
-Formally, during the $t$-th phase of training, the dataset $\mathcal{D}^{t}=(\mathcal{P},\mathcal{Y}^{t})=\{(p_{i},y_{i}^{t})\}_{i=1}^{N}$ is defined, where $\mathcal{P}$ represents the point clouds shared across all phases, and $\mathcal{Y}^{t}$ contains annotations exclusively for objects belonging to the class subset $\mathcal{C}^{t}$.
-It is important to note that the 3D scenes within each phase can contain objects of any type from the entire object category set $\mathcal{C}$, but only the object belonging to $\mathcal{C}^{t}$ are annotated during that specific phase.
-After training for phase $t$ completes, the model is evaluated on a validation set containing the union of classes up to task $t$ (i.e., $\mathcal{C}^{1}\cup\cdots\cup\mathcal{C}^{t}$).
-Training progresses to the next phase, $t+1$, where the model $\Phi$ observes the same set of 3D scenes $\mathcal{P}$ but with annotations for different object types belonging to the subset $\mathcal{C}^{t+1}$.
-This incremental training approach allows the model to gradually learn and adapt to new object categories over multiple phases.
-4.2 CLIMB-3D
-In our proposed framework (Figure 2), the incremental instance segmentation model undergoes phased training, as described in Sec.
-4.1, where carefully designed subsets of the dataset are introduced to handle various real-world scenarios discussed in Sec.
-4.3.
-Formally, at phase $t$, when the model $\Phi^{t}$ is introduced with input data $\mathcal{D}^{t}=\{(p_{i},y_{i}^{t})\}_{i=1}^{N}$ and trained using Eq.
-1, a common issue arises where it tends to forget the knowledge acquired in the previous phase, leading to catastrophic forgetting [mccloskey1989catastrophic].
-To address this, we first draw inspiration from techniques developed in the 2D domain [rebuffi2017icarl, li2017learning] and recent 3D semantic segmentation [su2024balanced, Yang_2023_CVPR], and adapt them for our setting.
-However, we observe that these adaptations alone fall short of achieving the desired performance levels; therefore, we propose a teacher-student knowledge distillation framework to effectively retain previously learned knowledge.
-Additionally, we incorporate an imbalance correction module to handle the challenge of less frequent classes from earlier tasks.
+3DIS.
+The objective of this task is to detect and segment each object instance within a given point cloud.
+Formally, the training dataset is denoted by $\mathcal{D}=\{(\mathbf{P}_{i},\mathbf{Y}_{i})\}_{i=1}^{N}$, where $N$ is the number of training samples.
+Each input $\mathbf{P}_{i}\in\mathbb{R}^{M\times 6}$ is a colored point cloud consisting of $M$ points, where each point is represented by its 3D coordinates and RGB values.
+The corresponding annotation $\mathbf{Y}_{i}=\{(m_{i,j},c_{i,j})\}_{j=1}^{J_{i}}$ contains $J_{i}$ object instances, where $m_{i,j}\in\{0,1\}^{M}$ is a binary mask indicating the points belonging to the $j$-th instance, and $c_{i,j}\in\mathcal{C}=\{1,\dots,C\}$ is the semantic class label of that instance.
+Given $\mathbf{P}_{i}$, the model $\Phi$ predicts instance-level outputs $\hat{\mathbf{Y}}_{i}=\{(\hat{m}_{i,j},\hat{c}_{i,j})\}_{j=1}^{\hat{J}_{i}}$, where $\hat{m}_{i,j}$ and $\hat{c}_{i,j}$ denote the predicted mask and category label for the $j$-th instance, respectively.
+The number of predicted instances $\hat{J}_{i}$ varies depending on the model’s inference.
+The model is optimized using below objective:
+$\mathcal{L}_{\text{3DIS}}(\mathcal{D};\Phi)$ $=\frac{1}{|\mathcal{D}|}\sum_{(\mathbf{P},\mathbf{Y})\in\mathcal{D}}\frac{1}{|\mathbf{Y}|}\sum_{(m_{j},c_{j})\in\mathbf{Y}}$
+$\quad\left(\mathcal{L}_{\text{mask}}(m_{j},\hat{m}_{j})+\lambda\cdot\mathcal{L}_{\text{cls}}(c_{j},\hat{c}_{j})\right)$ (1)
+where $\mathcal{L}_{\text{cls}}$ and $\mathcal{L}_{\text{mask}}$ are the average classification and mask losses over all instances, and $\lambda$ controls their trade-off.
+CI-3DIS.
+Unlike conventional 3DIS, CI-3DIS involves sequential learning, where object categories are introduced incrementally over $T$ training tasks.
+Each task $t$ introduces a disjoint set of classes $\mathcal{C}^{t}$, with $\mathcal{C}=\bigcup_{t=1}^{T}\mathcal{C}^{t}$ and $\mathcal{C}^{t}\cap\mathcal{C}^{t^{\prime}}=\emptyset$ for $t\neq t^{\prime}$.
+At each task $t$, the model receives a dataset $\mathcal{D}^{t}=\{(\mathbf{P}_{i},\mathbf{Y}_{i}^{t})\}_{i=1}^{N}$, where each coloured point cloud $\mathbf{P}_{i}\in\mathbb{R}^{M\times 6}$ contains $M$ points, and the corresponding annotation $\mathbf{Y}_{i}^{t}=\{(m_{i,j}^{t},c_{i,j}^{t})\}_{j=1}^{J_{i}^{t}}$ includes instance masks $m_{i,j}^{t}\in\{0,1\}^{M}$ and semantic labels $c_{i,j}^{t}\in\mathcal{C}^{t}$.
+The model $\Phi^{t}$, initialized from $\Phi^{t-1}$, is trained on $\mathcal{D}^{t}$ to predict instance-level outputs $\hat{\mathbf{Y}}_{i}^{t}=\{(\hat{m}_{i,j}^{t},\hat{c}_{i,j}^{t})\}_{j=1}^{\hat{J}_{i}^{t}}$, where each predicted class label $\hat{c}_{i,j}^{t}$ belongs to the cumulative set of all classes observed so far: $\bigcup_{k=1}^{t}\mathcal{C}^{k}$.
+The optimisation objective for the CI-3DIS setting is defined as:
+$\mathcal{L}_{\text{CI-3DIS}}(\mathcal{D}^{t};\Phi^{t})$ $=\frac{1}{|\mathcal{D}^{t}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in\mathcal{D}^{t}}\frac{1}{|\mathbf{Y}|}\sum_{(m^{t}_{j},c^{t}_{j})\in\mathbf{Y}^{t}}$
+$\quad\left(\mathcal{L}_{\text{mask}}(m^{t}_{j},\hat{m}^{t}_{j})+\lambda\cdot\mathcal{L}_{\text{cls}}(c^{t}_{j},\hat{c}^{t}_{j})\right).$ (2)
+The key challenge in this setup is to learn new classes without forgetting those encountered in previous tasks, despite limited supervision and the absence of past labels.
+3.2 Method: CLIMB-3D
+Overview.
+As illustrated in Fig.
+2 and introduced in Sec.
+3.1, the proposed framework for CI-3DIS follows a phase-wise training strategy.
+In each phase, the model is exposed to a carefully selected subset of the dataset, designed to simulate real-world scenarios discussed in Sec.
+3.3.
+Naïvely training on such phased data leads to catastrophic forgetting [mccloskey1989catastrophic], where the model forgets previously acquired knowledge when learning new tasks.
+To address this, we first incorporate Exemplar Replay (ER) [buzzega2020dark], inspired by 2D [rebuffi2017icarl, li2017learning, aljundi2018memory] and 3D [boudjoghra20243d] incremental setting, to mitigate forgetting by storing a small set of representative samples from earlier phases.
+However, ER alone is insufficient for achieving robust performance.
+We therefore introduce two additional components: a Pseudo-Label Generator (PLG), which leverages a frozen model to generate supervision signals for previously seen classes, and a Class-Balanced Re-weighting (CBR) module that compensates for class imbalance across phases.
+Each of these components is described in detail below  [^2] For clarity, modifications introduced by each component – ER, PLG, and CBR – are highlighted in green, blue, and red, respectively..
 Exemplar Replay (ER).
-Inspired by the approach proposed by Buzzega et al.
-[buzzega2020dark], ER methods alleviate the issue of limited exposure to previous task data during training.
-By selectively storing a small subset of exemplars $\mathcal{E}_{t}$ from previous phases, the model can learn from both the current task data $\mathcal{D}_{t}$ and the replayed exemplars $\mathcal{E}_{1:t-1}$.
-This results in a combined dataset $\mathcal{D}_{t}\cup\mathcal{E}_{1:t-1}$, where $\mathcal{E}_{1:t-1}$ represents the exemplar memory formed by the union of all previous exemplar sets $\mathcal{E}_{1:t-1}=\mathcal{E}_{1}\cup~{}\dots~{}\cup\mathcal{E}_{t-1}$.
-The model undergoes a full iteration on $\mathcal{D}_{t}$ before replaying the exemplars.
-Some previous 3D approaches adopt this strategy to retain knowledge but rely on a large exemplar set [boudjoghra20243d], which is often impractical in real-world scenarios.
-To address this, we choose a smaller exemplar set, creating a more challenging setup that requires the model to effectively manage and retain knowledge with limited resources, thereby testing its robustness and adaptability in practical applications.
-Knowledge Distillation (KD) Module.
-In our incremental learning approach, we utilize a Knowledge Distillation (KD) module that incorporates a teacher-student framework, maintaining a copy of the previously trained model while learning the current task.
-For $t>1$, at the beginning of each training stage, the current model $\Phi^{t}$ is initialized as $\Phi^{t}\leftarrow\Phi^{t-1}$, where $\Phi^{t-1}$ represents the model trained in the previous phase.
-As the $\Phi^{t}$ is trained on the previous stage dataset, it holds information about the previous set of classes.
-Hence, we make use of this model to help retain previous knowledge while learning the current task.
-When presented with a new training point-cloud and label pair $(p,y^{t})$, the output of the previous model is calculated as $\hat{y}_{\mathrm{pseudo}}=\Phi^{\mathrm{t-1}}(p)$, and a combined loss function is minimized.
-This combined loss comprises the $\mathcal{L}_{\mathrm{Seg}}(\hat{y}^{t},y^{t})$ loss, which measures the discrepancy between the predicted and ground truth labels, and the knowledge distillation loss $\mathcal{L}_{\mathrm{KD}}$, which encourages the similarity between the predictions of the current model and the previous model.
-$\mathcal{L}_{\mathrm{KD}}(\hat{y}^{t},\hat{y}_{\mathrm{pseudo}})$ $=\mathcal{L}_{\mathrm{mask}}(\hat{m}^{t}_{j},\hat{m}_{j,\mathrm{pseudo}})$
-$\quad\quad\quad\quad+\lambda_{\mathrm{cls}}\mathcal{L}_{\mathrm{cls}}(\hat{c}^{t}_{j},\hat{c}_{j,\mathrm{pseudo}})$ (2)
-However, as pointed out in previous works on object detection in 2D [liu2023continual], Section 4.2 is often biased towards the background classes, as the model tends to predict the background for most instances.
-Similarly, we propose selecting the top $K$ most confident predictions from the previous teacher model $\Phi^{t-1}$ and combining them with the ground truth labels, which then serve as pseudo-labels.
-By extracting the top $K$ confident samples from the output of $\Phi^{t-1}$ and combining them with the ground truth labels, the augmented label set becomes $y^{\prime}=[y^{t},\hat{y}_{\mathrm{pseudo}}^{K}]$.
-This augmented label set is then used to optimize the current model $\Phi$ with the $\mathcal{L}_{\mathrm{Seg}}$ loss function from Equation 1.
-Imbalance Correction (IC) Module.
-Although retaining a few samples from previous tasks and selecting the most confident predictions from the previous model, $\Phi^{t-1}$, helps preserve information from prior tasks while learning new ones, we observed that this approach does not adequately address the class imbalance.
-Our analysis reveals that the most confident predictions from $\Phi^{t-1}$ are largely associated with the most frequent object categories, causing the model to forget less common classes.
-This issue can be mitigated by re-weighting the predictions based on the frequency of observed categories [islam2021class].
-At task $t$, we only have access to the data and statistics of the current task; the previous task’s dataset and statistics are unavailable.
-To incorporate balancing elements despite the absence of previous stage statistics, we propose leveraging pseudo prediction statistics from $\Phi^{t-1}$.
-During each iteration, we use $\Phi^{t-1}$ to generate pseudo labels and accumulate class frequency statistics for prior tasks throughout the current epoch training.
-At the end of each epoch, we combine the statistics of observed scene classes and predicted pseudo-classes, calculating the frequency $\mathbf{f}$ of all classes seen so far.
-Formally, for each category $c$, we assign a weight $\mathbf{w}_{c}$ inversely proportional to its observed frequency in the predictions of $\Phi^{t-1}$ and the current dataset.
-The weight $\mathbf{w}_{c}$ is defined as: $\mathbf{w}_{c}=\frac{1}{\mathbf{f}(c)+\epsilon}$, where $\epsilon$ is a small constant to avoid division by zero.
-In the next epoch, predictions from $\Phi^{t-1}$ are re-weighted using $\mathbf{w}_{c}$, creating an adjusted high-confidence pseudo label set for less frequent categories: $\hat{y}_{\mathrm{pseudo}}^{(w)}=\mathbf{w}_{c}\cdot\hat{y}_{\mathrm{pseudo}}$.
-This re-weighting occurs at each epoch, allowing $\Phi^{t-1}$ to yield a broader set of less frequent classes.
-To ensure the model encounters both high-confidence and less frequent classes, we select the top $K$ high-confidence predictions both before and after re-weighting.
-The resulting augmented label space, which combines ground truth labels, original pseudo labels, and re-weighted pseudo labels, is given by: $y^{\prime}=[y,\hat{y}_{\mathrm{pseudo}}^{K},y_{\mathrm{pseudo}}^{(w),K}]$.
-As this weighting scheme is applied only to previous model predictions, we further tune the current model to favor less frequent classes by incorporating the same weights $\mathbf{w}_{c}$ into the classification loss of Equation 1.
-The adjusted segmentation loss becomes:
-$\mathcal{L}_{\mathrm{3DIS}}(y^{\prime}_{j},\hat{y}_{j})=\mathcal{L}_{\mathrm{mask}}(m^{\prime}_{j},\hat{m}_{j})+\mathbf{w}^{\prime}_{c}\mathcal{L}_{\mathrm{cls}}(c^{\prime}_{j},\hat{c}_{j}),$ (3)
-where $\mathbf{w}^{\prime}_{c}=\mathbf{w}_{c}\cdot\lambda_{\mathrm{cls}}$ represents the adjusted category weights.
-By re-weighting and augmenting both the label space and the loss function, our IC module ensures that both the current and previous models encounter pseudo labels spanning the long-tail distribution, addressing class imbalance and enabling more balanced learning.
-4.3 Designing Incremental Scenarios
-Figure 3: Incremental scenarios are grouped based on frequency of occurrence, semantic similarity, and random clustering.
-Different color clouds ( ,  ,  ) represent tasks in each scenario, while various shapes represent object categories, and   denotes the background.
-Left: Tasks are organized based on the frequency of object categories.
-Middle: Tasks are grouped by semantic similarity, where objects with similar shapes (e.g., circles, plus signs, and triangles) denote semantically similar classes.
-Right: In this fully random scenario, tasks may contain a mix of semantically similar, more frequent, or less frequent classes.
-While conventional incremental learning methods have numerous practical applications, they often assume an equal distribution of samples, which does not reflect real-world conditions.
-In practice, the number of object categories, $\mathcal{C}$, is typically large, with significant variability in category occurrence, shape, structure, and size.
-With these attributes in mind, we design three incremental learning scenarios, each addressing distinct aspects of real-world conditions, the design is highlighted in Figure 3.
-1
-Frequency Scenarios (Split_A).
-This scenario acknowledges that datasets are often labeled based on the frequency of category occurrences.
+Inspired by Buzzega et al.
+[buzzega2020dark], ER addresses the issue of catastrophic forgetting by allowing the model to retain a small portion of data from earlier phases.
+At phase $t$, the model is trained not only on the current task data $\mathcal{D}^{t}$ but also on a set of exemplars $\mathcal{E}^{1:t-1}$ collected from previous phases.
+These exemplars are accumulated incrementally, where $\mathcal{E}^{1:t-1}=\mathcal{E}^{1}\cup\dots\cup\mathcal{E}^{t-1},$ and $\mathcal{E}^{t}$ denotes the exemplar subset stored after phase $t$, with $|\mathcal{E}^{t}|$ as the exemplar replay size.
+We define the combined replay dataset as: ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}=\mathcal{D}^{t}\cup\mathcal{E}^{1:t-1}.$ Epoch training proceeds in two stages: first on the current data $\mathcal{D}^{t}$, followed by replay using $\mathcal{E}^{1:t-1}$.
+The updated objective function for CI-3DIS, based on LABEL:eq:ci3dis-loss, becomes:
+$\mathcal{L}_{\text{CI-3DIS}}({\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}};\Phi^{t})$ $=\frac{1}{|{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}}\frac{1}{|\mathbf{Y}^{t}|}\sum_{(m_{j}^{t},c_{j}^{t})\in\mathbf{Y}^{t}}$
+$\quad\left(\mathcal{L}_{\text{mask}}(m_{j}^{t},\hat{m}_{j}^{t})+\lambda\cdot\mathcal{L}_{\text{cls}}(c_{j}^{t},\hat{c}_{j}^{t})\right).$ (3)
+While previous CI-3DIS approaches use large exemplar sets to mitigate forgetting [boudjoghra20243d], such strategies become impractical under memory constraints.
+Consequently, although exemplar replay (ER) aids knowledge retention, it alone proves insufficient for robust performance with a limited exemplar budget.
+We therefore propose PLG and CBM to better preserve past representations and enhance generalisation across phases.
+Pseudo-Label Generator (PLG).
+In the CI-3DIS setting, at phase $t$ (for $t>1$), although we no longer have access to the ground-truth labels for previously seen classes, the model from the previous phase, $\Phi^{t-1}$, remains available and retains knowledge acquired from past tasks.
+We use this model to generate pseudo-labels for previous classes, thereby providing approximate past supervision during training.
+This allows the current model $\Phi^{t}$ to retain past knowledge and reduces forgetting, even without access to original annotations.
+Given a point cloud and label pair $(\mathbf{P},\mathbf{Y}^{t})$ from $\mathcal{D}^{t}$, the previous model model $\Phi^{t-1}$ generates pseudo-labels for the previously learned classes as $\hat{\mathbf{Y}}^{1:t-1}=\Phi^{t-1}(\mathbf{P})$, which we denote as $\tilde{\mathbf{Y}}^{t}$ for brevity.
+However, we observe that using all predictions from $\Phi^{t-1}$ introduces noisy or incorrect labels, which degrade overall performance.
+To mitigate this, we select the top-$K$ most confident instance predictions, denoted as $\tilde{\mathbf{Y}}^{t}_{K}$, and concatenate them with the ground-truth labels of the current task to obtain the supervision set: ${\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{\mathbf{Y}}^{t}}=\mathbf{Y}^{t}\mathbin{\|}{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\tilde{\mathbf{Y}}^{t}_{K}}$.
+At the end, Using this supervision set, the model is optimized via the following updated objective:
+$\mathcal{L}_{\text{CI-3DIS}}({\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}};\Phi^{t})$ $=\frac{1}{|{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}}\frac{1}{|{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{\mathbf{Y}}^{t}}|}\sum_{({\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{m}_{j}^{t}},{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{c}_{j}^{t}})\in{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{\mathbf{Y}}^{t}}}$
+$\quad\left(\mathcal{L}_{\text{mask}}({\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{m}_{j}^{t}},\hat{m}_{j}^{t})+\lambda\cdot\mathcal{L}_{\text{cls}}({\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{c}_{j}^{t}},\hat{c}_{j}^{t})\right),$ (4)
+where the loss is computed over both the current task’s ground-truth labels and the top-$K$ confident pseudo-labels from previous phases, enabling the model to retain prior knowledge while adapting to new classes.
+Class-Balanced Re-weighting (CBR).
+While ER and PLG contribute to preserving prior knowledge, we observe that the model’s predictions are biased toward frequently occurring classes, leading to forgetting of rare categories.
+This issue is further reinforced by the top-$K$ pseudo-labels selected by PLG, which predominantly represent dominant classes in the data.
+To alleviate this issue, we introduce a Class-Balanced Re-weighting (CBR) scheme that adjusts for class imbalance based on frequency of object occurrence.
+However, at task $t$, we only have access to the dataset $\mathcal{D}^{t}$ and its corresponding label distribution $p^{t}(c)$ over classes $c\in\mathcal{C}^{t}$.
+The datasets and class distributions from previous tasks $\{\mathcal{D}^{i},p^{i}(c)\}_{i=1}^{t-1}$ are no longer accessible – making it challenging to directly account for class imbalance across all seen tasks.
+Therefore, we propose to leverage the pseudo-label predictions of the frozen model $\Phi^{t-1}$ as a proxy for the class distribution across previously learned categories.
+During each training iteration, we apply the previous model $\Phi^{t-1}$ to the current input point cloud $\mathbf{P}$ to generate pseudo-labels $\tilde{\mathbf{Y}}^{t}$ , from which we accumulate class-wise frequency statistics corresponding to prior tasks.
+Let $\mathcal{C}^{1:t}=\bigcup_{i=1}^{t}\mathcal{C}^{i}$ represent the union of all classes seen up to task $t$.
+At the end of each epoch, we compute the overall class frequency $\mathbf{f}\in\mathbb{R}^{|\mathcal{C}^{1:t}|}$ by combining the pseudo-label distribution $\tilde{p}^{t}(c)$ derived from $\tilde{\mathbf{Y}}^{t}$ with the observed ground-truth label distribution $p^{t}(c)$ from the current task.
+The resulting frequency vector $\mathbf{f}=[f(c)]_{c\in\mathcal{C}^{1:t}}$ is then used to compute class-wise $\mathbf{w}=[w(c)]_{c\in\mathcal{C}^{1:t}}$ weights for balanced pseudo-label selection and to tune current model to favour rare classes.
+During the next epoch, the model $\Phi^{t-1}$ produces soft pseudo-label predictions $\tilde{\mathbf{Y}}^{t}$ , where each component $\tilde{\mathbf{Y}}^{t}[c]$ denotes the confidence score for class $c\in\mathcal{C}^{1:t-1}$.
+To promote balanced pseudo-label selection, we apply class-wise re-weighting using $w(c)=\frac{1}{f(c)+\epsilon}$, where $\epsilon$ is a small constant for numerical stability.
+The adjusted predictions are computed by applying the class-wise weights as: $\tilde{\mathbf{Y}}^{t,w}[c]=w(c)\cdot\tilde{\mathbf{Y}}^{t}[c],\quad\forall c\in\mathcal{C}^{1:t-1}.$ We then perform top-$K$ selection over both the original and the re-weighted scores, yielding two pseudo-label sets: $\tilde{\mathbf{Y}}^{t}_{K}$ and $\tilde{\mathbf{Y}}^{t,w}_{K}$, respectively.
+The final supervision label set $\overline{\mathbf{Y}}^{t}$ is constructed by concatenating the ground-truth labels $\mathbf{Y}^{t}$ with both sets of pseudo-labels: ${\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}=\mathbf{Y}^{t}\mathbin{\|}{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\tilde{\mathbf{Y}}^{t}_{K}}\mathbin{\|}{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\tilde{\mathbf{Y}}^{t,w}_{K}}.$
+Although the above re-weighting mitigates bias from prior tasks, it does not address class imbalance within the current task $t$.
+To resolve this, we extend the class-balanced weights $w_{c}$ to incorporate statistics from both previously seen and current classes, i.e., over $\mathcal{C}^{1:t}$ rather than only $\mathcal{C}^{1:t-1}$, and denote the updated weights as $w_{c}^{\prime}$, which are then integrated into the training objective.
+After incorporating all components, the modified objective in Eq.
+1 becomes the final updated objective used in CLIMB-3D:
+$\mathcal{L}_{\text{CI-3DIS}}({\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}};\Phi^{t})$ $=\frac{1}{|{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}}\frac{1}{|{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}|}\sum_{({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{m}_{j}^{t}},{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{c}_{j}^{t}})\in{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}}$
+$\quad\left(\mathcal{L}_{\text{mask}}({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{m}_{j}^{t}},\hat{m}_{j}^{t})+{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime\prime}_{c}}\cdot\mathcal{L}_{\text{cls}}({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{c}_{j}^{t}},\hat{c}_{j}^{t})\right),$ (5)
+where ${\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime\prime}_{c}}={\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime}_{c}}\cdot\lambda_{\mathrm{cls}}$ denotes the scaled weight vector, and $\lambda_{\mathrm{cls}}$ is a loss-balancing hyperparameter.
+To address the class-imbalance across tasks, our CBR module jointly re-weights pseudo-label selection and classification loss, ensuring balanced supervision across all seen categories, including rare ones.
+This unified strategy reduces bias and enhances performance, particularly for under-represented classes.
+3.3 Benchmarking Incremental Scenarios
+While CI-3DIS methods offer a wide range of practical applications, they frequently rely on the assumption of uniform sample distribution – an assumption that rarely holds in real-world settings.
+In practice, the number of object categories, denoted by $\mathcal{C}$, is often large and characterised by substantial variability in category frequency, shape, structure, and size.
+To address these challenges, we propose three incremental learning scenarios, each designed to capture a different facet of real-world complexity.
+The design of these scenarios is detailed below; for more information and illustration, please refer Appendix Appendix C of the supplementary material.
+Split_A: Frequency Scenarios.
+This scenario acknowledges that datasets are often labelled based on the frequency of category occurrences.
 To accommodate this, we propose a split where the model initially learns from the most frequent categories and subsequently incorporates the less frequent ones in later stages.
-By prioritizing the training of frequently occurring categories, the model can establish a strong foundation before expanding its knowledge to handle rarer categories.
-2
-Semantic Scenarios (Split_B).
-In real-world environments, objects may exhibit similarities in appearance, and then moved to different environments, the model may encounter new objects that does not share similar semantic characteristics with previously seen categories.
-To address this, we introduce the Split_B  scenario.
-It involves grouping categories based on their semantic labels and incrementally training the model to handle these groups.
-This allows the model to generalize its knowledge across semantically similar categories, facilitating adaptation to new objects with similar characteristics.
-Unlike the Split_A  scenario, this scenario may include both frequent and infrequent categories within the same task.
-3
-Random Scenarios (Split_C).
-In some cases, data labeling is based on the availability of objects rather than specific criteria.
-To account for this scenario, we design the Split_C  scenario.
-This scenario represents a completely random setting where each task can have any class, leading to varying degrees of class imbalance.
-By exposing the model to such diverse and imbalanced scenarios, we aim to enhance its ability to handle real-world situations where the availability of labeled data is unpredictable.
-By designing these three incremental learning scenarios, we aim to provide a more realistic representation of object distributions, frequencies, and dynamics encountered in the real world.
+By prioritizing the training of frequently occurring categories, the model can establish a strong foundation before expanding its knowledge to handle rare categories.
+Split_B: Semantic Scenarios.
+Beyond frequency, semantic similarity is crucial in real-world deployments.
+While objects often share visual or functional traits, models may encounter semantically different categories in new environment.
+To simulate this challenge, we introduce the Split_Bscenario.
+Here, categories are grouped based on their semantic relationships, and the model is incrementally trained on one semantic group at a time.
+This setup encourages the model to generalise across semantically similar categories and adapt more effectively when exposed to new ones.
+Unlike the Split_Ascenario, which organises learning based on category frequency, the Split_Bscenario may contain both frequent and infrequent categories within a single task, focusing instead on semantic continuity and transfer.
+Split_C: Random Scenarios.
+In some cases, data labelling is driven by the availability of objects rather than predefined criteria.
+To capture this, we introduce the Split_Cscenario, which represents a fully random setting where any class can appear in any task, resulting in varying levels of class imbalance.
+By exposing the model to such diverse and unpredictable distributions, we aim to improve its robustness in real-world situations where labelled data availability is inconsistent.
+By designing these incremental scenarios, we aim to provide a more realistic representation of object distributions, frequencies, and dynamics encountered in the real world.
 4 Experiments
-We evaluate our method across three scenarios specifically designed to capture the complexities of real-world settings, where new categories emerge incrementally alongside class imbalance, as discussed in Section 4.3.
-Experimental results demonstrate that our approach effectively handles the gradual introduction of new classes and mitigates the impact of class imbalance within these scenarios.
-The following sections detail the datasets, evaluation metrics, incremental scenarios, and implementation procedures, followed by a thorough analysis and comparison of the results.
-5.1 Experimental Setup
-Datasets.
-We evaluate our method using the ScanNet200 dataset [rozenberszki2022language], which includes 200 object categories and exhibits inherent class imbalance, making it well-suited for simulating and evaluating real-world scenarios.
-Additionally, we benchmark our method against existing incremental learning approaches using the original ScanNet dataset [dai2017scannet] in a semantic segmentation setting.
-We follow the standard training and validation splits as defined in prior works.
-Evaluation Metrics.
-We evaluate our method using mean Average Precision (mAP), a standard metric for 3D instance segmentation that provides a comprehensive measure of segmentation quality, accounting for both precision and recall.
-For comparison with existing semantic incremental learning approaches, we report mean Intersection over Union (mIoU), which measures the overlap between predicted and ground truth instances, offering a detailed evaluation of segmentation accuracy.
-To assess the model’s ability to mitigate catastrophic forgetting in continual learning scenarios, we use the Forgetting Percentage Points (FPP), as defined in [liu2023continual].
-This metric quantifies performance degradation by measuring the accuracy drop between the initial and final training phases, on the categories observed in the first training phase.
-Incremental Scenarios.
-As discussed in Section 4.3, we design three incremental scenarios: Split_A, Split_B, and Split_C, each consisting of three tasks, which are grouped based on object occurrence frequency, semantic similarity, and random grouping, respectively.
-In Split_A, the frequency of object categories progressively decreases through the tasks.
-This scenario follows the head, common, and tail splits present in the ScanNet200 dataset, with class distributions of 66-68-66 in each split.
-In Split_B, we partition the classes into 74-50-76 based on semantic similarity, which is calculated using the CLIP [radford2021learning] text encoder, followed by clustering using K-Means.
-Finally, in Split_C, the classes are shuffled and split into three sets, resulting in 67-67-66 categories per split.
-These scenarios allow for a comprehensive evaluation of our approach under varying conditions, facilitating a deeper understanding of its performance and generalization in diverse real-world settings.
-Implementation Details.
-We utilize the transformer-based model for 3D instance segmentation proposed in [Schult23ICRA], designed to iteratively attend to hierarchical feature representations.
-The model processes the 4 coarsest levels of a ResNet-based U-Net backbone across three iterations, progressively refining from coarse to fine, resulting in $L=12$ transformer decoder layers.
-Each transformer decoder layer shares weights across iterations and consists of a standard transformer layer utilizing self-attention and masked cross-attention mechanisms.
-The feature backbone employed is Minkowski Res16UNet34C [choy20194d].
-Training Details.
-We adopt the data augmentation, hyperparameters, and training strategy described in [Schult23ICRA].
-For joint training (Row 1, Table 4), the model is trained for 600 epochs using the AdamW optimizer [loshchilov2017decoupled] with a one-cycle learning rate scheduler [smith2018superconvergence], and results are evaluated on the entire validation set.
-In incremental training, we retain the same hyperparameters, adjusting only the number of epochs, and we use a memory buffer size of 50 scenes.
-Training is divided into three phases, introducing one split per phase across the three designed scenarios.
-After each phase, the resulting model is evaluated on all classes encountered up to that point.
-5.2 Results and Discussion
-Table 1: Comparison between the baseline and proposed method with mAP25, mAP50, and mAP, which is after training for all the stages. We also report the FPP metric.
+Table 1: Comparison between the proposed method and the baseline in the 3DIS setting, evaluated using mAP25, mAP50, mAP, and FPP after training across all phase.
 Scenarios Methods Average Precision $\uparrow$ FPP $\downarrow$
@@
 Baseline 16.46 14.29 10.44 51.30 46.82
-Split_A CLIMB-3D 35.69 31.05 22.72 3.44 2.63
+Split_A CLIMB-3D (Ours) 35.69 31.05 22.72 3.44 2.63
 Baseline 17.22 15.07 10.93 46.27 42.1
-Split_B CLIMB-3D 35.48 31.56 23.69 8.00 5.51
+Split_B CLIMB-3D (Ours) 35.48 31.56 23.69 8.00 5.51
 Baseline 25.65 21.08 14.85 31.68 28.84
-Split_C CLIMB-3D 31.59 26.78 18.93 9.10 7.89
-To evaluate our proposed method, we conduct a comparative analysis using exemplar replay (ER) for instance segmentation and [Yang_2023_CVPR] for semantic segmentation as baselines.
-As shown in Table 1, our method, which integrates exemplar replay, knowledge distillation, and an imbalance correction (IC) module, achieves notable improvements over the baseline in terms of mAP and FPP.
-Specifically, in the Split_A  scenario, our approach significantly enhances overall performance.
-We observe an improvement of 19.23%, 31.05%, and 12.28% for $\text{mAP}_{50}$, $\text{mAP}_{25}$, and overall mAP, respectively, while reducing forgetting by 47.86% as measured by $\text{mAP}_{50}$.
-For the Split_B  scenario as well, CLIMB-3D demonstrates a consistent performance boost over the baseline, significantly improving mAP and reducing forgetting, which is lowered to 5.52% compared to 46.21% in the baseline for AP50.
-Likewise, in the Split_C  scenario, CLIMB-3D enhances both learning efficiency and forgetting reduction, achieving a performance of 26.78% in $\text{mAP}_{50}$  and reducing forgetting by 20.95% compared to baseline.
-These results across scenarios underscore the effectiveness of our approach.
-Table 2: Comparison with previous method on semantic segmentation on ScanNet V2 dataset. We report the mIoU metric for the evaluated models.
+Split_C CLIMB-3D (Ours) 31.59 26.78 18.93 9.10 7.89
+Table 2: Comparison between the proposed method and existing baselines in the 3DSS setting on ScanNetV2 [dai2017scannet], evaluated using mIoU.
 Methods Phase=1 Phase=2 All
@@
 LwF [li2017learning] 30.38 13.37 26.13
-Yang et al. [Yang_2023_CVPR] 34.16 13.43 28.98
+Yang Yang_2023_CVPR 34.16 13.43 28.98
 CLIMB-3D (Ours) 69.39 32.56 59.38
-Although our method focuses on segmenting individual objects (instance segmentation), we also demonstrate its performance in semantic segmentation by presenting a comparative analysis with existing methods for class-incremental semantic segmentation on the ScanNet V2 dataset (Table 2).
-Using our predicted labels, we assign each point the label corresponding to the highest confidence mask and exclude background labels (floor and wall), as these are not part of the object-level segmentation.
-Following the dataset splits established by [Yang_2023_CVPR], we report results for both training phases.
-Our proposed method achieves a substantial improvement over prior methods, with a gain of 35.23% in Phase 1 and approximately 19.1% in Phase 2.
-Overall, our method reaches a mIoU of 59.38%, significantly outperforming previous baselines, which achieve a lower mIoU of around 30%.
+Table 3: Ablation study showing the impact of each component ER, PLG, and CBR in a three-phase setup.
+Each split (s) and number corresponds to data introduced at that phase.
+The All column reports performance over all classes seen so far.
+Joint training (Oracle) is highlighted in gray, and the full model in orange.
+Best results are in bold.
+Row Modules p=1 $\uparrow$ p=2 $\uparrow$ p=3 $\uparrow$ FPP $\downarrow$
+s1 s1 s2 All s1 s2 s3 All
+1. Oracle - - - - 55.14 30.77 25.30 37.68 -
+2. Naïve 56.82 0.00 28.09 14.15 0.00 0.00 19.67 5.80 56.82
+3. + ER 56.82 18.51 32.81 25.72 10.38 9.43 24.27 14.28 46.44
+4. + PLG 56.82 50.00 34.39 42.13 49.78 11.41 26.47 29.28 7.04
+5. + CBR 56.82 54.67 33.75 44.13 54.19 12.02 26.55 31.05 2.63
+4.1 Setup
+Datasets.
+We evaluate our method on the ScanNet200 [rozenberszki2022language], which comprises 200 object categories and exhibits significant class imbalance – making it ideal for simulating and assessing real-world scenarios.
+In addition, we benchmark our approach against existing incremental learning methods using the ScanNetV2 [dai2017scannet] in a 3DSS setting.
+For this evaluation, we follow the standard training and validation splits defined in prior works [Yang_2023_CVPR].
+Evaluation Metrics.
+We evaluate our method using mean Average Precision (mAP), a standard metric for 3DIS that provides a comprehensive measure of segmentation quality by accounting for both precision and recall.
+For comparison with existing 3DSS methods, we report the mean Intersection over Union (mIoU), which quantifies the overlap between predicted and ground truth segments.
+To assess the model’s ability to mitigate catastrophic forgetting in incremental settings, we use the Forgetting Percentage Points (FPP) metric, as defined in [liu2023continual].
+FPP captures performance degradation by measuring the accuracy drop on the initially seen categories between the first and final training phases.
+Incremental Scenarios.
+As discussed in Section 3.3, we design three incremental scenarios – Split_A, Split_B, and Split_C– each with three tasks, based on object frequency, semantic similarity, and random grouping, respectively.
+For Split_A, we follow the class split provided by rozenberszki2022language, with a class distribution of 66-68-66 across the three tasks.
+In Split_B, we partition the classes into 74-50-76 based on semantic similarity, which is calculated using the CLIP [radford2021learning] text encoder, followed by clustering using K-Means.
+Finally, in Split_C, classes are randomly shuffled and split into three sets, with 67-67-66 categories per split.
+These scenarios comprehensively evaluate our approach’s performance and generalization in varied real-world conditions.
+Implementation Details.
+For all the experiments we use Mask3D [Schult23ICRA] with the Minkowski Res16UNet34C [choy20194d] as a feature backbone.
+We adopt the data augmentation, hyperparameters, and training strategy described in [Schult23ICRA], with using the AdamW optimizer [loshchilov2017decoupled] and a one-cycle learning rate scheduler [smith2018superconvergence], and results are evaluated on the entire validation set.
+In incremental training, learning is divided into three phases, introducing one split per phase across the three scenarios described in Sec.
+3.3.
+For each phase, we train for 500, 300, and 300 epochs, respectively, except for the Oracle setting (Row 1 in Tab.
+3), which is trained once for 600 epochs.
+We retain the same hyperparameters throughout, modifying only the number of training epochs.
+The exemplar replay size is fixed at $|\mathcal{E}^{t}|=50$ scenes.
+4.2 Results and Discussion
+To evaluate our proposed approach, we construct a baseline using ER [buzzega2020dark] for the 3DIS setting, as no existing incremental baselines are available for this task.
+In contrast, for the 3DSS setting, where prior incremental baselines do exist, we compare our method against three existing approaches [serra2018overcoming, Yang_2023_CVPR, li2017learning].
+Results on CI-3DIS.
+Tab.
+1 compares CLIMB-3D against the ER baseline across three CI-3DIS scenarios, evaluated after all phases using $\text{mAP}_{25}$, $\text{mAP}_{50}$, overall mAP, and FPP.
+In the Split_Ascenario, characterized by significant distribution shifts, it achieves gains of +19.23%, +16.76%, and +12.28% in $\text{mAP}_{25}$, $\text{mAP}_{50}$, and overall mAP, while drastically reducing forgetting with FPP scores of 3.44% and 2.63%, over 47 points lower than the baseline.
+Under the Split_Bscenario, with semantically related new categories, improvements of +18.26%, +16.49%, and +12.76% are observed alongside a reduction in forgetting to 8.00% and 5.51% from baseline levels above 42%.
+For the Split_Cscenario, featuring increasing geometric complexity, it maintains solid gains of +5.94%, +5.70%, and +4.08% across mAP metrics, with forgetting reduced by over 22 points.
+Overall, these results demonstrate robust forward transfer, minimal forgetting, and stable learning, validating the method’s effectiveness for scalable continual 3D semantic understanding.
+Results on CI-3DSS.
+Although our primary focus is on CI-3DIS, we additionally evaluate our method under the CI-3DSS setting to enable comparison with existing approaches.
+For this, we adapt our predictions by assigning each point the label corresponding to the highest-confidence mask and exclude background classes (floor and wall), as these are not part of the semantic segmentation targets.
+Tab.
+2 presents a comparison between our method and existing baselines on the ScanNet V2 dataset [dai2017scannet], evaluated using mIoU across two training phases and overall.
+Despite being designed for instance segmentation, our method generalises effectively to semantic segmentation, achieving substantial gains: +35.23% mIoU in Phase 1 and +19.1% in Phase 2.
+Overall, it achieves 59.38% mIoU – significantly higher than the $\sim$30mIoU of prior methods – highlighting its robustness and transferability.
+For a detailed per-phase analysis, evaluation of rare classes, and qualitative result comparisons, please refer to Appendix Appendix A, Appendix Appendix B, and Appendix Appendix D, respectively, in the supplementary material.
+4.3 Ablation
+We conduct an ablation study to evaluate the contribution of each component in our framework.
+As an upper bound, we report the performance of an Oracle model trained jointly on the full dataset.
+For the incremental setting, we follow the previously defined splits and first train the model naively across phases.
+We then incrementally add each module to isolate its impact.
+Tab.
+3 presents results on the Split_Ascenario using both $\text{mAP}_{50}$  and FPP metrics.
+Naïve Training.
+When trained naïvely without any dedicated modules, the model suffers from severe catastrophic forgetting, as evident in row 2.
+The model completely forgets previously learned classes upon entering new phases, with performance dropping to zero for earlier splits and a high FPP of 56.82.
+Effect of ER.
+Adding exemplar replay (row 3) partially alleviates forgetting by maintaining a buffer of past examples.
+This yields notable gains for s1 in phase 2 (+18.51%) and phase 3 (+10.38%), and also improves s2 in phase 3 (+9.43%).
+However, the overall forgetting remains substantial (FPP: 46.44), showing ER alone is insufficient.
+Effect of PLG.
+The addition of the pseudo-label generator (PLG), which generates labels for previous classes by retaining a copy of the model from earlier phase, facilitates knowledge retention and forward transfer.
+As shown in row 4, PLG significantly reduces forgetting and enhances performance on current tasks.
+For s1, it improves $\text{mAP}_{50}$ by 31.49% in phase 2 and 39.40% in phase 3 over exemplar replay.
+Overall, PLG yields a 15.00% increase in performance and reduces forgetting by 39.40%.
+Effect of CBR.
+Finally, class-balanced re-weighting (CBR) mitigates class imbalance both during pseudo-labelling and current task learning by adjusting each class’s contribution based on its frequency.
+As shown in row 5, CBR enhances s1 retention over PLG in phases 2 and 3 (+4.67% and +4.41%) and further improves performance on s2 and s3.
+It also achieves the lowest FPP of 2.63, reflecting effective forgetting mitigation.
+Overall, CBR offers the best trade-off between retaining past knowledge and learning new tasks, outperforming PLG and ER by 4.41% and 43.81%, respectively.
+5 Conclusion
+We address the critical challenge of catastrophic forgetting in class-incremental 3D instance segmentation by proposing a modular framework that combines exemplar replay (ER), a pseudo-label generator (PLG), and class-balanced re-weighting (CBR).
+Our method incrementally adapts to new classes while preserving knowledge from previous tasks without requiring access to the full dataset.
+Through extensive experiments on a three-phase benchmark, we demonstrate the individual and combined effectiveness of each component.
+ER provides a easy knowledge storage, PLG enhances knowledge retention via pseudo-label supervision, and CBR further refines learning by mitigating class imbalance.
+Together, these modules significantly reduce forgetting and improve overall segmentation performance across all phases.
+By balancing stability and plasticity, our method contributes to continual 3D scene understanding and establishing a new baseline for future research.
+Supplementary Material
 Table 4: Comparison of results in terms of mAP50 with and proposed CLIMB-3D for three different scenarios.
@@
 Split_C CLIMB-3D 36.40 32.63 33.38 33.00 28.51 17.11 34.64 26.78
-We extend the analysis from Table 1 to Table 3 to highlight the impact of our proposed method on individual splits across various scenarios.
-The results clearly demonstrate that our model consistently retains knowledge of previous tasks better than the baseline.
-For Split_A , our model shows improvement throughout the phase.
-In Phase 3 of (s2), although both the baseline and our method exhibit a performance drop, our method reduces forgetting significantly compared to the baseline.
-The Split_B  scenario, while more complex than Split_A, achieves comparable results due to semantic similarity among classes within the same task.
-In Phase 2, our model achieves overall all 43.13% $\text{mAP}_{50}$ compared to 24.53% on baseline, a similar trend is observed in Phase 3, where our method not only consistently improves learning but also enhances retention of previous information.
-After all three tasks, our method achieves an overall performance of 31.56% AP50, compared to 15.07% for the baseline.
-In the Split_C  scenario, the first-stage model struggles due to the increased complexity introduced by random grouping.
-In Phase 2, while the baseline focuses on learning the current task, it suffers from severe forgetting of prior knowledge.
-Conversely, our method balances new task learning with the retention of earlier information.
-By Phase 3, the model effectively consolidates s1 and maintains strong performance across all task splits.
-Overall, our proposed method improves mAP by 5.6%.
-5.3 Ablation
-To assess the effectiveness of each component in our proposed framework, we perform an ablation study.
-Initially, we establish an upper-bound performance by jointly training the model on the complete dataset using a transformer-based architecture, such as Mask3D [Schult23ICRA], referred to as the Oracle.
-For the incremental learning setup, we generate training splits according to the scenarios outlined earlier.
-In this study, we first train the model naively across phases and then sequentially integrate each module to evaluate its individual contribution to performance.
-Table 4 summarizes the results for the Split_A  scenario, using both the $\text{mAP}_{50}$  and FPP metrics.
-Table 4: Ablation study results illustrating the impact of exemplar replay, knowledge distillation, and imbalance correction modules in a three-phase training setup.
-Each split, representing the subset of data introduced at each phase (p), is labeled as ‘s’ followed by the phase number.
-The final column, ‘All’, in each phase reports performance across all classes encountered up to that phase.
-Joint training results (Oracle) are highlighted in gray, while results with all modules combined are marked in orange.
-The best-performing results are shown in bold.
-Row Modules p=1 $\uparrow$ p=2 $\uparrow$ p=3 $\uparrow$ FPP $\downarrow$
-s1 s1 s2 All s1 s2 s3 All
-1. Oracle - - - - 55.14 30.77 25.30 37.68 -
-2. Naïve 56.82 0.00 28.09 14.15 0.00 0.00 19.67 5.80 56.82
-3. + ER 56.82 18.51 32.81 25.72 10.38 9.43 24.27 14.28 46.44
-4. + KD 56.82 50.00 34.39 42.13 49.78 11.41 26.47 29.28 7.04
-5. + IC 56.82 54.67 33.75 44.13 54.19 12.02 26.55 31.05 2.63
-Naïve Training.
-In the naive incremental training setup, where no dedicated modules are incorporated, the model learns the current task but suffers from catastrophic forgetting of the previously learned tasks, as expected.
-This behavior is evident in row 2, where, upon transitioning to phase 2, the model entirely forgets the classes learned during phase 1.
-A similar trend is observed in phase 3, and this pattern is also reflected in the FPP metric.
-Effect of Exemplar Replay.
-To mitigate catastrophic forgetting, we incorporate exemplar replay, which stores and replays examples from previous tasks.
-As shown in row 3, exemplar replay improves average precision by 18.5% for s1 in phase 2 and 10.38% in phase 3.
-It also reduces forgetting for s2 by 9.43% in phase 3, while slightly improving learning on the current task.
-However, substantial forgetting persists, as reflected in the FPP metric, highlighting the limitations of exemplar replay alone.
-Effect of Knowledge Distillation.
-The addition of knowledge distillation (KD), which retains a copy of the model from previous tasks, facilitates the preservation of past task knowledge while enabling forward knowledge transfer.
-As shown in row 4, KD considerably reduces forgetting and boosts performance on the current task.
-Specifically, for s1, KD improves $\text{mAP}_{50}$ by 31.49% in phase 2 and by 39.40% in phase 3, compared to exemplar replay.
-Overall, KD leads to a 15.00% increase in performance while reducing forgetting by 39.40% after all tasks have been learned.
-Effect of Imbalance Correction.
-The imbalance correction module addresses the class imbalance in the dataset by re-weighting the teacher model’s predictions during KD based on class frequency.
-As highlighted in row 5 of Table 4, this addition further improves performance.
-Specifically, for s1, imbalance correction reduces forgetting by 4.67% and 4.41% in phases 2 and 3, respectively, compared to the results without this module (row 4).
-For s2, while a slight decrease in current task performance is observed in phase 2, this is likely due to the module’s prioritization of mitigating forgetting less frequent classes in previous tasks.
-In phase 3, performance on s2 and s3 improves.
-Overall, imbalance correction significantly reduces forgetting, achieving improvements of 4.41% and 43.81% over KD and exemplar replay, respectively.
-6 Conclusion
-We address the challenge of class-incremental 3D instance segmentation with class imbalance.
-We propose an innovative approach that integrates a memory-efficient exemplar replay buffer, knowledge distillation, and a novel imbalance correction module.
-This framework mitigates the forgetting of rare classes during incremental learning by accounting for the frequency of object occurrences.
-To enable comprehensive evaluation, we design three incremental learning scenarios, each comprising three phases that reflect real-world dynamics.
-Our experimental results demonstrate that the proposed framework significantly enhances the learning of new classes while reducing forgetting of previously learned ones.
-The carefully designed scenarios and framework not only offer a strong baseline but also provide a clear benchmark for future research, laying a foundation for more advanced techniques in class-incremental learning.
-Supplementary Material
-In this supplementary material, we first demonstrate the performance gains on rare classes achieved by incorporating the IC module in Appendix A.
-Next, we provide detailed split information for all scenarios, based on class names, in Appendix B.
-Finally, we present a qualitative comparison between the baseline method and our proposed approach in Appendix C.
-Appendix Appendix A Evaluation on Rare Categories
-The proposed imbalance correction (IC) module, as detailed in Section 4.2, is designed to address the performance gap for rare classes.
-To assess its impact, we compare its performance with the framework which has exemplar replay (ER) and knowledge distillation (KD).
-Specifically, we focus on its ability to improve performance for rare classes, which the model encounters infrequently compared to more common classes.
 Table 5: Results for classes observed by the model 1–20 times during an epoch, evaluated on Split_A  for Phase 2, in terms of $\text{mAP}_{50}$.
@@
 Average - 46.66 54.98
-The results, shown in Table 5 and Table 6, correspond to evaluations on Split_A  for Phase 2 and Phase 3, respectively.
-In Phase 2, we evaluate classes seen 1–20 times per epoch, while Phase 3 targets even less frequent classes, with observations limited to 1–10 times per epoch.
-As illustrated in Table 5, the IC module substantially improves performance on rare classes in terms of $\text{mAP}_{50}$  in Phase 2 of Split_A.
-For instance, classes like recycling bin and trash bin, seen only 3 and 7 times, respectively, shows significant improvement when the IC module is applied.
-Overall, the IC module provides an average boost of 8.32%, highlighting its effectiveness in mitigating class imbalance.
 Table 6: Results for classes observed by the model 1–10 times during an epoch, evaluated on Split_A  for Phase 3, in terms of $\text{mAP}_{50}$.
@@
 Average - 32.08 44.21
-Similarly, Table 6 presents results for Phase 3, demonstrating significant gains for infrequent classes.
-For example, even though the classes such as piano, bucket, and laundry basket are observed only once, IC module improves the performance by 52.30%, 10.40%, and 13.60%, respectively.
-The ER+KD module does not focus on rare classes like shower and toaster which results in low performance, but the IC module compensates for this imbalance by focusing on underrepresented categories.
-On average, the addition of the proposed IC module into the framework outperforms ER+KD by 12.13%.
+Figure 3: Tasks are grouped into incremental scenarios based on object frequency, semantic similarity, and random assignment.
+,  , and   denote different tasks; shapes indicate object categories;   marks the background.
+Left: Grouped by category frequency.
+Middle: Grouped by semantic similarity (e.g., similar shapes).
+Right: Randomly grouped, mixing semantic and frequency variations.
 Table 7: Classes grouped by tasks for each proposed scenario on the ScanNet200 dataset labels.
@@
 potted plant
-Appendix Appendix B Incremental Scenarios Phases
-Table 7 presents the task splits for each proposed scenario introduced in Section 4.3 using the ScanNet200 dataset.
-The three scenarios, Split_A, Split_B, and Split_C, are each divided into three tasks: Task 1, Task 2, and Task 3.
-Notably, the order of classes in these tasks is random.
-Appendix Appendix C Qualitative Results
+In this supplementary material, we first present the per-phase performance of the proposed CLIMB-3D model in Appendix Appendix A, followed by the analysis of gains on rare classes achieved through the incorporation of the CBR module in Appendix Appendix B.
+Next, we provide detailed split information for all scenarios, based on class names, in Appendix Appendix C.
+Finally, we present a qualitative comparison between the baseline method and our proposed approach in Appendix Appendix D.
+Appendix Appendix A Per-Phase Performance Analysis
+We extend the analysis presented in Tab.
+1 (Sec.
+4.2 of the main paper) to Tab.
+4 to examine the impact of our proposed method across individual splits and phases in different scenarios.
+The results consistently show that our method outperforms the baseline by better retaining knowledge from earlier tasks.
+Split_A.
+In the Split_Ascenario, our model shows notable improvements across all phases.
+Particularly in Phase 3, although both methods show a drop in performance for s2, our approach significantly reduces forgetting, maintaining a higher $\text{mAP}_{50}$  across all splits.
+Split_B.
+The Split_Bscenario, while more challenging due to increased semantic overlap, demonstrates our model’s strength in generalisation.
+In Phase 2, it achieves an overall $\text{mAP}_{50}$  of 43.13% compared to the baseline’s 24.53%.
+This trend continues in Phase 3, with our model achieving 31.56%, more than double the baseline’s 15.07%, indicating both improved learning and retention.
+Split_C.
+For the Split_Cscenario, where splits are randomly grouped, Phase 1 performance is initially lower.
+However, in subsequent phases, our method balances stability and plasticity better than the baseline.
+By Phase 3, it consolidates performance across all splits, improving the overall $\text{mAP}_{50}$to 26.78%, a 5.7% gain over the baseline.
+These results demonstrate the robustness of our method in handling continual learning under varied and challenging split strategies.
+Appendix Appendix B Evaluation on Rare Categories
+The proposed Class-Balanced Re-weighting (CBR) module, as detailed in 3.2, is designed to address the performance gap for rare classes.
+To assess its impact, we compare its performance with the framework which has exemplar replay (ER) and Pseudo-Label Generator (PLG).
+Specifically, we focus on its ability to improve performance for rare classes, which the model encounters infrequently compared to more common classes.
+The results, shown in Tab.
+5 and Table 6, correspond to evaluations on Split_A  for Phase 2 and Phase 3, respectively.
+In Phase 2, we evaluate classes seen 1–20 times per epoch, while Phase 3 targets even less frequent classes, with observations limited to 1–10 times per epoch.
+As illustrated in Table 5, the CBR module substantially improves performance on rare classes in terms of $\text{mAP}_{50}$  in Phase 2 of Split_A.
+For instance, classes like recycling bin and trash bin, seen only 3 and 7 times, respectively, show significant improvement when the CBR module is applied.
+Overall, the CBR module provides an average boost of 8.32%, highlighting its effectiveness in mitigating class imbalance.
+Similarly, Table 6 presents results for Phase 3, demonstrating significant gains for infrequent classes.
+For example, even though classes such as piano, bucket, and laundry basket are observed only once, the CBR module improves the performance by 52.30%, 10.40%, and 13.60%, respectively.
+The ER+PLG module does not focus on rare classes like shower and toaster, which results in low performance, but the CBR module compensates for this imbalance by focusing on underrepresented categories.
+On average, the addition of the proposed CBR module into the framework outperforms ER+PLG by 12.13%.
+Appendix Appendix C Incremental Scenarios Phases
+Table 7 presents the task splits for each proposed scenario introduced in Sec.
+3.3, and Fig.
+3 illustrates them visually for the ScanNet200 [rozenberszki2022language] dataset.
+The three scenarios – Split_A, Split_B, and Split_C– are each divided into three tasks: Task1, Task2, and Task3.
+Notably, the order of classes within these tasks is random.
+Appendix Appendix D Qualitative Results
 In this section, we present a qualitative comparison of the proposed framework with the baseline method.

### v3

@@
-CLIMB-3D: Continual Learning for Imbalanced 3D Instance Segmentation
+
 Abstract
-While 3D instance segmentation (3DIS) has advanced significantly, existing methods typically assume that all object classes are known in advance and are uniformly distributed.
+While 3D instance segmentation (3DIS) has advanced significantly, most existing methods assume that all object classes are known in advance and uniformly distributed.
 However, this assumption is unrealistic in dynamic, real-world environments where new classes emerge gradually and exhibit natural imbalance.
-Although some approaches have addressed class emergence, they often overlook class imbalance, resulting in suboptimal performance – particularly on rare categories.
-To tackle this challenge, we propose CLIMB-3D, a unified framework for CLass-incremental Imbalance-aware 3DIS.
-Building upon established exemplar replay (ER) strategies, we show that ER alone is insufficient to achieve robust performance under constrained memory conditions.
-To mitigate this, we introduce a novel pseudo-label generator (PLG) that extends supervision to previously learned categories by leveraging predictions from a frozen prior model.
-Despite its promise, PLG tends to bias towards frequent classes.
-Therefore, we propose a class-balanced re-weighting (CBR) scheme, that estimates object frequencies from pseudo-labels and dynamically adjusts training bias – without requiring access to past data.
-We design and evaluate three incremental scenarios for 3DIS on the challenging ScanNet200 dataset, and additionally on semantic segmentation on ScanNetV2.
-Our approach achieves state-of-the-art results, surpassing prior work by up to 16.76% mAP for instance segmentation and approximately 30% mIoU for semantic segmentation, demonstrating strong generalization across both frequent and rare classes.
+Although some approaches address the emergence of new classes, they often overlook class imbalance, which leads to suboptimal performance, particularly on rare categories.
+To tackle this, we propose CLIMB-3D, a unified framework for CLass-incremental Imbalance-aware 3DIS.
+Building upon established exemplar replay (ER) strategies, we show that ER alone is insufficient to achieve robust performance under memory constraints.
+To mitigate this, we introduce a novel pseudo-label generator (PLG) that extends supervision to previously learned categories by leveraging predictions from a frozen model trained on prior tasks.
+Despite its promise, PLG tends to be biased towards frequent classes.
+Therefore, we propose a class-balanced re-weighting (CBR) scheme that estimates object frequencies from pseudo-labels and dynamically adjusts training bias, without requiring access to past data.
+We design and evaluate three incremental scenarios for 3DIS on the challenging ScanNet200 dataset and additionally validate our method for semantic segmentation on ScanNetV2.
+Our approach achieves state-of-the-art results, surpassing prior work by up to 16.76% mAP for instance segmentation and approximately 30% mIoU for semantic segmentation, demonstrating strong generalisation across both frequent and rare classes.
+Code is available at: https://github.com/vgthengane/CLIMB3D.
 1 Introduction
-Figure 1: Overview of the CI-3DIS setting: new classes are introduced incrementally as tasks progress.
-After each task, the model must recognize both previously learned and newly added classes.
-For instance, at Task 2, classes like Pillow, Coffee Table, and Sofa Chair are introduced, while the model retains the ability to detect earlier classes such as Table, Chair, and Couch.
-3D instance segmentation (3DIS) is an essential task in computer vision with widespread applications in graphics, robotics, and augmented reality [boudjoghra20243d, lai2022stratified].
-It involves identifying and segmenting individual objects in 3D space, enabling precise object boundaries and class labels that enhance scene understanding and perception in dynamic environments.
-In recent years, a variety of methods for 3DIS have been proposed, including top-down [wang2018sgpn, jia2021scaling, zhang2021point], bottom-up [hou20193dsis, yang2019learning], and transformer-based approaches [Schult23ICRA].
-These methods have demonstrated strong performance under the traditional setting, which assumes that all object classes are available and well-balanced during training.
-However, these assumptions limit their applicability in real world settings, where new categories gradually emerge over time, often exhibiting natural class imbalance.
-These limitations in current 3DIS methods underscore the need for a Class-Incremental Learning (CIL) [rebuffi2017icarl, de2021continual] framework capable of learning new categories as they appear, while maintaining accuracy on previously learned classes – especially rare ones, which are more susceptible to without catastrophic forgetting [mccloskey1989catastrophic].
-Unfortunately, most existing research in CIL has focused on the 2D domain for image classification [rebuffi2017icarl, li2017learning, aljundi2018memory, serra2018overcoming], with limited extensions to object detection [liu2023continual, joseph2021towards, shmelkov2017incremental] and semantic segmentation [cermelli2020modeling, cermelli2022incremental, douillard2021plop].
-Extensions to 3D point clouds remain limited, with most work focusing on object-level classification rather than scene-level segmentation [dong2021i3dol, liu2021l3doc, chowdhury2021learning].
-Recent works on scene-level class-incremental 3DIS (CI-3DIS) [boudjoghra20243d] and class-incremental semantic segmentation (CI-3DSS) [Yang_2023_CVPR] show promise, but typically rely on large exemplar buffers [rolnick2019experience] Moreover, they typically overlook class imbalance, limiting real-world applicability.
+Figure 1: Overview of the CI-3DIS setting.
+New object categories are introduced incrementally with each task.
+After every phase, the model must recognise both newly added and previously learned classes.
+For instance, in Task 2, categories such as Pillow, Coffee Table, and Sofa Chair are introduced, while the model is expected to retain recognition of earlier classes like Table, Chair, and Couch.
+3D instance segmentation (3DIS) is a fundamental task in computer vision that involves identifying precise object boundaries and class labels in 3D space, with broad applications in graphics, robotics, and augmented reality [boudjoghra20243d, lai2022stratified].
+Traditional 3DIS methods, including top-down [wang2018sgpn, jia2021scaling, zhang2021point], bottom-up [hou20193dsis, yang2019learning], and transformer-based approaches [Schult23ICRA], perform well under the assumption that all classes are available and balanced during training.
+However, this assumption does not hold in real-world environments, where new object classes emerge over time and exhibit natural imbalance.
+This gap motivates the need for Class-Incremental Learning (CIL) [rebuffi2017icarl, de2021continual], which supports learning new categories while retaining knowledge of previous ones [mccloskey1989catastrophic].
+While CIL has seen success in 2D image tasks [rebuffi2017icarl, li2017learning, aljundi2018memory, serra2018overcoming], extensions to 3D point clouds remain limited, often focusing on object-level classification [dong2021i3dol, liu2021l3doc, chowdhury2021learning].
+Recent works on scene-level class-incremental 3DIS (CI-3DIS) [boudjoghra20243d] and class-incremental semantic segmentation (CI-3DSS) [Yang_2023_CVPR] show promise, but depend heavily on large exemplar [rolnick2019experience] and overlook class imbalance, limiting their practicality.
 To address this, we propose CLIMB-3D, a unified framework for CLass-incremental IMBalance-aware 3DIS that jointly tackles catastrophic forgetting and class imbalance.
 The framework begins with Exemplar Replay (ER), storing a small number of samples from past classes for later replay.
-However, this alone does not yield promising results under the strict memory constraints needed CI-3DIS.
-To address this, we introduce a Pseudo-Label Generator (PLG), which uses a frozen previous task model to generate supervision for earlier classes.
-However, we observed that, PLG tends to favour frequent classes while ignoring under-represented ones.
+However, this alone does not yield promising results under the strict memory constraints required in CI-3DIS.
+To address this, we introduce a Pseudo-Label Generator (PLG), which uses a frozen model from previous task to generate supervision for earlier classes.
+However, we observed that PLG tends to favour frequent classes while ignoring under-represented ones.
 To mitigate this, we propose a Class-Balanced Re-weighting (CBR), which estimates class frequencies from the pseudo-labels to enhance the learning of rare classes without accessing past data.
 Together, these components form a practical and effective solution for real-world 3DIS.
-To evaluate CLIMB-3D in a realistic incremental setup, we design three benchmark scenarios using the ScanNet200 [rozenberszki2022language] dataset, which contains 200 classes with a natural long-tail distribution.
-These scenarios simulate real-world conditions where new classes gradually emerge, following inherently imbalanced distributions.
-Specifically, new classes appear in groups based on one of three criteria: (A) object frequency, (B) semantic similarity, or (C) random grouping without a specific pattern.
-Our experiments on these scenarios for incremental 3DIS show that CLIMB-3D effectively mitigates forgetting of previous tasks while improving performance across both frequent and rare classes, delivering consistent gains over prior methods.
-In summary, our contributions are threefold: (i) We introduce a novel problem setting of imbalanced class-incremental 3DIS, along with a simple and effective method to mitigate catastrophic forgetting and ensure balanced learning; (ii) We establish three benchmark scenarios that realistically model continuous object emergence with natural class imbalance; and (iii) We validate our framework’s effectiveness, achieving SOTA results with improvements up to 16.76% mAP over existing baselines [^1] Code will be available at: https://github.com/vgthengane/CLIMB3D .
+To evaluate CLIMB-3D, we designed three benchmark scenarios for CI-3DIS based on the ScanNet200 dataset [rozenberszki2022language].
+These scenarios simulate incremental learning setup under natural class imbalance, where new classes emerge based on (A) object frequency, (B) semantic similarity, or (C) random grouping.
+Additionally, for comparison with existing methods, we evaluate our approach on CI-3DSS using the ScanNetV2 dataset [dai2017scannet].
+Experimental results show that our method significantly reduces forgetting and improves performance across both frequent and rare categories, outperforming previous methods in both settings.
+Fig.˜1 illustrates the CI-3DIS setup.
+In summary, our contributions are: (i) a novel problem setting of imbalanced class-incremental 3DIS with an effective method to balance learning and mitigate forgetting; (ii) three benchmarks modelling continual object emergence with natural imbalance; and (iii) strong experimental results, achieving up to 16.76% mAP improvement over baselines.
 2 Related Work
-This section reviews the current literature on 3D instance segmentation and incremental learning methods, including the limited work addressing incremental learning for 3D scene-level tasks.
-2.1 3D Instance Segmentation
-Various approaches have been proposed for 3D instance segmentation.
-One common approach adopts a bottom-up pipeline, in which an embedding in the latent space is learned to facilitate the clustering of object points [wang2018sgpn, lahoud20193d, jiang2020pointgroup, zhang2021point, chen2021hierarchical, han2020occuseg, he2021dyco3d, liang2021instance].
-These methods are also known as grouping-based or clustering-based methods.
-Other methods use a top-down approach, also known as proposal-based methods, where 3D bounding boxes are first detected, then the object region is segmented within the box [yang2019learning, hou20193dsis, liu2020learning, yi2019gspn, engelmann20203d].
-Recently, the transformer architecture [vaswani2017attention] has also been used for the task of 3D instance segmentation [Schult23ICRA, sun2022superpoint], motivated by work in 2D [cheng2022masked, cheng2021per].
-While these methods propose various models for improving the quality of the object segments, they rely on the availability of annotations for all object categories.
-On the other hand, we target learning in a progressive manner, in which new semantic annotation is provided and past data is inaccessible.
-In order to reduce the annotation cost for 3D instance segmentation, various methods propose weakly supervised alternatives to methods that use dense annotations [xie2020pointcontrast, hou2021exploring, chibane2022box2mask].
-While these methods improve the ability to learn from a small set of annotated examples, they rely on a fixed set of semantic labels, so they are prone to catastrophic forgetting in an incremental setting.
-2.2 Incremental Learning
-Incremental, lifelong, or continual learning methods aim to train a machine learning model sequentially to avoid “catastrophic forgetting” which is caused by training the model on a set of data and later training on another set of data.
-There are several methods have been proposed for this paradigm, these methods can be divided into three categories: (i) Model Regularization [kirkpatrick2017overcoming, aljundi2018memory, li2017learning] methods limit the plasticity of model parameters to avoid catastrophic forgetting of previous tasks.
-These methods include weight regularization such as EWC [serra2018overcoming] and function regularization such as knowledge distillation [hinton2015distilling].
-(ii) Exemplar replay approaches either create a subset of the past task data or generate samples using generative models to avoid privacy concerns and save those in memory to replay while learning new tasks.
-This method is effective in more challenging settings and datasets [rebuffi2017icarl, kamra2017deep, buzzega2020dark, cha2021co2l].
-(iii) Dynamic network expansion-based method learns a new task by either dynamically expanding the model [rusu2016progressive, li2019learn, zhao2022deep] or by creating a subset of the model [zhao2022deep, wang2020learn, ke2020continual, rajasegaran2019random] to learn to cater for a new task.
-Recent approaches to 3D class-incremental segmentation, such as [Yang_2023_CVPR] and [su2024balanced], have made some initial contributions.
-However, these methods often fall short in performance as they do not leverage state-of-the-art 3D segmentation models and are primarily focused on semantic segmentation, while our work emphasizes object-level instance segmentation.
-Kontogianni et al.
-[kontogianni2024continual] propose a general online continual learning framework and evaluate it on 3D dataset segmentation.
-Similarly, [boudjoghra20243d] addresses the open-world 3D incremental learning problem but relies heavily on an extensive memory buffer.
-In contrast, our work introduces a dedicated continual learning framework for 3D instance segmentation, with a focus on effective knowledge transfer from previous tasks, while also accounting for the challenges posed by infrequent class occurrences.
+3D Instance Segmentation.
+Various methods have been proposed for 3DIS.
+Grouping-based approaches adopt a bottom-up pipeline that learns latent embeddings for point clustering [wang2018sgpn, lahoud20193d, jiang2020pointgroup, zhang2021point, chen2021hierarchical, han2020occuseg, he2021dyco3d, liang2021instance].
+In contrast, proposal-based methods follow a top-down strategy by detecting 3D boxes and segmenting objects within them [yang2019learning, hou20193dsis, liu2020learning, yi2019gspn, engelmann20203d].
+Recently, transformer-based models [vaswani2017attention] have also been applied to 3DIS and 3DSS [Schult23ICRA, sun2022superpoint], inspired by advances in 2D vision [cheng2022masked, cheng2021per].
+However, these approaches require full annotations for all classes and are not designed for progressive learning, where only new class annotations are available and previous data is inaccessible.
+Another line of work aims to reduce annotation costs by proposing weakly supervised 3DIS methods based on sparse cues [xie2020pointcontrast, hou2021exploring, chibane2022box2mask].
+While effective with limited annotations, these methods assume a fixed set of classes and are susceptible to catastrophic forgetting in incremental settings.
+Incremental Learning.
+Continual learning involves training models sequentially to mitigate catastrophic forgetting.
+One common strategy is model regularisation, which constrains parameter updates using techniques such as elastic weight consolidation or knowledge distillation [kirkpatrick2017overcoming, aljundi2018memory, li2017learning, serra2018overcoming, hinton2015distilling].
+Another approach is exemplar replay, where past data is stored or generated to rehearse older tasks while learning new ones [rebuffi2017icarl, kamra2017deep, buzzega2020dark, cha2021co2l].
+A third direction involves dynamically expanding the model architecture or using modular sub-networks to accommodate new tasks without interfering with old ones [rusu2016progressive, li2019learn, zhao2022deep, wang2020learn, ke2020continual, rajasegaran2019random].
+Recent efforts in class-incremental 3D segmentation [Yang_2023_CVPR, su2024balanced] have shown early promise but often rely on basic architectures and focus primarily on semantic segmentation.
+Other works [kontogianni2024continual, boudjoghra20243d] explore continual learning in 3D settings, though they often depend on large memory buffers.
+In contrast, our work introduces a tailored framework for 3D instance segmentation that effectively transfers knowledge across tasks, even under class imbalance.
+Long-tailed Recognition.
+Imbalanced class distributions lead to poor recognition of rare (long-tail) classes.
+To address this, re-sampling methods balance the data via input [chawla2002smote, gupta2019lvis, park2022majority, shen2016relay, mahajan2018exploring] or feature space [ando2017deep, chu2020feature, li2021metasaug], avoiding naïve under-/over-sampling.
+Loss re-weighting offers another direction, using class-based [khan2017cost, he2022relieving, cao2019learning, khan2019striking, cui2019class, wang2021adaptive] or per-example [lin2017focal, ren2018learning, shu2019meta] adjustments to ensure fair contribution.
+Parameter regularisation improves generalisation through weight constraints [alshammari2022long], albeit requiring careful tuning.
+Other approaches leverage transfer learning [yin2019feature, zhou2020bbn], self-supervision [yang2020rethinking, li2021self], or contrastive learning [khosla2020supervised, li2021self, zhu2022balanced] to improve rare-class representations.
+Long-tailed recognition is well-studied in 2D with large-scale datasets [cui2019class, van2018inaturalist, liu2019large]; in 3D, ScanNet200 [rozenberszki2022language, dai2017scannet] enables similar efforts.
+Prior 3D work focused on re-weighting, re-sampling, and transfer learning [rozenberszki2022language], while regularisation remains underexplored.
+CeCo [zhong2023understanding] balances class centres via auxiliary loss but overlooks sampling and augmentation.
+Lahoud et al.
+[lahoud2024long] propose adaptive classifier regularisation for 3D segmentation, outperforming prior approaches without threshold tuning.
+However, neither method considers incremental learning.
+In contrast, we jointly address long-tail and continual 3D semantic segmentation.
 3 Methodology
 Figure 2: Overview of CLIMB-3D for CI-3DIS.
-The model incrementally learns new classes in sequential phases using ER to mitigate forgetting, a PLG for supervision on old classes, and CBR to handle class imbalance.
+The model incrementally learns new classes across sequential phases.
+During task $t$, point clouds $\mathbf{P}$ and their corresponding labels $\mathbf{Y}^{t}$ are sampled from a combination of the current training dataset $D^{t}$ and Exemplar Replay (ER), which maintains a small memory of past examples.
+These are then passed to the current model $\Phi^{t}$ to produce predicted labels $\mathbf{Y}^{t}$.
+The Pseudo-Label Generator (PLG) selects the top-$K$ predictions from the previous model $\Phi^{t-1}$.
+These pseudo-labels are then weighted based on class frequency $f(c)$ using Class-Balanced Re-weighting (CBR), and the top-$K$ re-weighted labels are selected to form the balanced pseudo-label set $\bar{\mathbf{Y}}^{t}$.
+This pseudo-label set is then concatenated with the ground-truth labels to form a final augmented supervision set $\overline{\mathbf{Y}}^{t}$ for task $t$, which is used to optimise the model $\Phi^{t}$ using Eq.˜3.
 3.1 Problem Formulation
 3DIS.
-The objective of this task is to detect and segment each object instance within a given point cloud.
-Formally, the training dataset is denoted by $\mathcal{D}=\{(\mathbf{P}_{i},\mathbf{Y}_{i})\}_{i=1}^{N}$, where $N$ is the number of training samples.
-Each input $\mathbf{P}_{i}\in\mathbb{R}^{M\times 6}$ is a colored point cloud consisting of $M$ points, where each point is represented by its 3D coordinates and RGB values.
-The corresponding annotation $\mathbf{Y}_{i}=\{(m_{i,j},c_{i,j})\}_{j=1}^{J_{i}}$ contains $J_{i}$ object instances, where $m_{i,j}\in\{0,1\}^{M}$ is a binary mask indicating the points belonging to the $j$-th instance, and $c_{i,j}\in\mathcal{C}=\{1,\dots,C\}$ is the semantic class label of that instance.
+The objective of this task is to detect and segment each object instance within a point cloud.
+Formally, the training dataset is denoted as $\mathcal{D}=\{(\mathbf{P}_{i},\mathbf{Y}_{i})\}_{i=1}^{N}$, where $N$ is the number of training samples.
+Each input $\mathbf{P}_{i}\in\mathbb{R}^{M\times 6}$ is a coloured point cloud consisting of $M$ points, where each point is represented by its 3D coordinates and RGB values.
+The corresponding annotation $\mathbf{Y}_{i}=\{(m_{i,j},c_{i,j})\}_{j=1}^{J_{i}}$ contains $J_{i}$ object instances, where $m_{i,j}\in\{0,1\}^{M}$ is a binary mask indicating which points belong to the $j$-th instance, and $c_{i,j}\in\mathcal{C}=\{1,\dots,C\}$ is the semantic class label of that instance.
 Given $\mathbf{P}_{i}$, the model $\Phi$ predicts instance-level outputs $\hat{\mathbf{Y}}_{i}=\{(\hat{m}_{i,j},\hat{c}_{i,j})\}_{j=1}^{\hat{J}_{i}}$, where $\hat{m}_{i,j}$ and $\hat{c}_{i,j}$ denote the predicted mask and category label for the $j$-th instance, respectively.
 The number of predicted instances $\hat{J}_{i}$ varies depending on the model’s inference.
-The model is optimized using below objective:
-$\mathcal{L}_{\text{3DIS}}(\mathcal{D};\Phi)$ $=\frac{1}{|\mathcal{D}|}\sum_{(\mathbf{P},\mathbf{Y})\in\mathcal{D}}\frac{1}{|\mathbf{Y}|}\sum_{(m_{j},c_{j})\in\mathbf{Y}}$
-$\quad\left(\mathcal{L}_{\text{mask}}(m_{j},\hat{m}_{j})+\lambda\cdot\mathcal{L}_{\text{cls}}(c_{j},\hat{c}_{j})\right)$ (1)
+The model is optimised using the following objective:
+$\mathcal{L}_{\text{3DIS}}(\mathcal{D};\Phi)=\frac{1}{|\mathcal{D}|}\sum_{(\mathbf{P},\mathbf{Y})\in\mathcal{D}}\frac{1}{|\mathbf{Y}|}\sum_{(m_{j},c_{j})\in\mathbf{Y}}\left(\mathcal{L}_{\text{mask}}(m_{j},\hat{m}_{j})+\lambda\cdot\mathcal{L}_{\text{cls}}(c_{j},\hat{c}_{j})\right),$ (1)
 where $\mathcal{L}_{\text{cls}}$ and $\mathcal{L}_{\text{mask}}$ are the average classification and mask losses over all instances, and $\lambda$ controls their trade-off.
@@
 Each task $t$ introduces a disjoint set of classes $\mathcal{C}^{t}$, with $\mathcal{C}=\bigcup_{t=1}^{T}\mathcal{C}^{t}$ and $\mathcal{C}^{t}\cap\mathcal{C}^{t^{\prime}}=\emptyset$ for $t\neq t^{\prime}$.
-At each task $t$, the model receives a dataset $\mathcal{D}^{t}=\{(\mathbf{P}_{i},\mathbf{Y}_{i}^{t})\}_{i=1}^{N}$, where each coloured point cloud $\mathbf{P}_{i}\in\mathbb{R}^{M\times 6}$ contains $M$ points, and the corresponding annotation $\mathbf{Y}_{i}^{t}=\{(m_{i,j}^{t},c_{i,j}^{t})\}_{j=1}^{J_{i}^{t}}$ includes instance masks $m_{i,j}^{t}\in\{0,1\}^{M}$ and semantic labels $c_{i,j}^{t}\in\mathcal{C}^{t}$.
-The model $\Phi^{t}$, initialized from $\Phi^{t-1}$, is trained on $\mathcal{D}^{t}$ to predict instance-level outputs $\hat{\mathbf{Y}}_{i}^{t}=\{(\hat{m}_{i,j}^{t},\hat{c}_{i,j}^{t})\}_{j=1}^{\hat{J}_{i}^{t}}$, where each predicted class label $\hat{c}_{i,j}^{t}$ belongs to the cumulative set of all classes observed so far: $\bigcup_{k=1}^{t}\mathcal{C}^{k}$.
-The optimisation objective for the CI-3DIS setting is defined as:
-$\mathcal{L}_{\text{CI-3DIS}}(\mathcal{D}^{t};\Phi^{t})$ $=\frac{1}{|\mathcal{D}^{t}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in\mathcal{D}^{t}}\frac{1}{|\mathbf{Y}|}\sum_{(m^{t}_{j},c^{t}_{j})\in\mathbf{Y}^{t}}$
-$\quad\left(\mathcal{L}_{\text{mask}}(m^{t}_{j},\hat{m}^{t}_{j})+\lambda\cdot\mathcal{L}_{\text{cls}}(c^{t}_{j},\hat{c}^{t}_{j})\right).$ (2)
+During each task $t$, the model receives a dataset $\mathcal{D}^{t}=\{(\mathbf{P}_{i},\mathbf{Y}_{i}^{t})\}_{i=1}^{N}$, where each coloured point cloud $\mathbf{P}_{i}\in\mathbb{R}^{M\times 6}$ contains $M$ points, and the corresponding annotation $\mathbf{Y}_{i}^{t}=\{(m_{i,j}^{t},c_{i,j}^{t})\}_{j=1}^{J_{i}^{t}}$ includes instance masks $m_{i,j}^{t}\in\{0,1\}^{M}$ and semantic class labels $c_{i,j}^{t}\in\mathcal{C}^{t}$.
+The model $\Phi^{t}$, initialized from $\Phi^{t-1}$, is trained on $\mathcal{D}^{t}$ to predict instance-level outputs $\hat{\mathbf{Y}}_{i}^{t}=\{(\hat{m}_{i,j}^{t},\hat{c}_{i,j}^{t})\}_{j=1}^{\hat{J}_{i}^{t}}$, where each predicted semantic label $\hat{c}_{i,j}^{t}$ belongs to the cumulative set of all classes observed so far: $\bigcup_{k=1}^{t}\mathcal{C}^{k}$.
 The key challenge in this setup is to learn new classes without forgetting those encountered in previous tasks, despite limited supervision and the absence of past labels.
@@
 Overview.
-As illustrated in Fig.
-2 and introduced in Sec.
-3.1, the proposed framework for CI-3DIS follows a phase-wise training strategy.
-In each phase, the model is exposed to a carefully selected subset of the dataset, designed to simulate real-world scenarios discussed in Sec.
-3.3.
-Naïvely training on such phased data leads to catastrophic forgetting [mccloskey1989catastrophic], where the model forgets previously acquired knowledge when learning new tasks.
-To address this, we first incorporate Exemplar Replay (ER) [buzzega2020dark], inspired by 2D [rebuffi2017icarl, li2017learning, aljundi2018memory] and 3D [boudjoghra20243d] incremental setting, to mitigate forgetting by storing a small set of representative samples from earlier phases.
-However, ER alone is insufficient for achieving robust performance.
-We therefore introduce two additional components: a Pseudo-Label Generator (PLG), which leverages a frozen model to generate supervision signals for previously seen classes, and a Class-Balanced Re-weighting (CBR) module that compensates for class imbalance across phases.
-Each of these components is described in detail below  [^2] For clarity, modifications introduced by each component – ER, PLG, and CBR – are highlighted in green, blue, and red, respectively..
+As shown in Fig.˜2 and formulated in Sec.˜3.1, the proposed framework for CI-3DIS follows a phase-wise training strategy.
+In each phase, the model is exposed to a carefully curated subset of the dataset, simulating real-world scenarios discussed in Sec.˜3.3.
+Training naïvely on such phased data leads to catastrophic forgetting [mccloskey1989catastrophic], where the model forgets prior knowledge when learning new tasks.
+To address this, we first incorporate Exemplar Replay (ER) [buzzega2020dark], a strategy inspired by 2D incremental methods [rebuffi2017icarl, li2017learning, aljundi2018memory] and recently extended to 3D settings [boudjoghra20243d], to mitigate forgetting by storing a small set of representative samples from earlier phases.
+However, ER alone is insufficient to achieve robust performance.
+To address this, two additional components are introduced: a Pseudo-Label Generator (PLG), which leverages a frozen model to generate supervision signals for previously seen classes, and a Class-Balanced Re-weighting (CBR) module that compensates for class imbalance across phases.
+Each of these components is described in detail below  [^1] For clarity, modifications introduced by each component (ER, PLG, and CBR) are highlighted in green, blue, and red, respectively..
 Exemplar Replay (ER).
 Inspired by Buzzega et al.
-[buzzega2020dark], ER addresses the issue of catastrophic forgetting by allowing the model to retain a small portion of data from earlier phases.
-At phase $t$, the model is trained not only on the current task data $\mathcal{D}^{t}$ but also on a set of exemplars $\mathcal{E}^{1:t-1}$ collected from previous phases.
-These exemplars are accumulated incrementally, where $\mathcal{E}^{1:t-1}=\mathcal{E}^{1}\cup\dots\cup\mathcal{E}^{t-1},$ and $\mathcal{E}^{t}$ denotes the exemplar subset stored after phase $t$, with $|\mathcal{E}^{t}|$ as the exemplar replay size.
-We define the combined replay dataset as: ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}=\mathcal{D}^{t}\cup\mathcal{E}^{1:t-1}.$ Epoch training proceeds in two stages: first on the current data $\mathcal{D}^{t}$, followed by replay using $\mathcal{E}^{1:t-1}$.
-The updated objective function for CI-3DIS, based on LABEL:eq:ci3dis-loss, becomes:
-$\mathcal{L}_{\text{CI-3DIS}}({\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}};\Phi^{t})$ $=\frac{1}{|{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}}\frac{1}{|\mathbf{Y}^{t}|}\sum_{(m_{j}^{t},c_{j}^{t})\in\mathbf{Y}^{t}}$
-$\quad\left(\mathcal{L}_{\text{mask}}(m_{j}^{t},\hat{m}_{j}^{t})+\lambda\cdot\mathcal{L}_{\text{cls}}(c_{j}^{t},\hat{c}_{j}^{t})\right).$ (3)
-While previous CI-3DIS approaches use large exemplar sets to mitigate forgetting [boudjoghra20243d], such strategies become impractical under memory constraints.
-Consequently, although exemplar replay (ER) aids knowledge retention, it alone proves insufficient for robust performance with a limited exemplar budget.
-We therefore propose PLG and CBM to better preserve past representations and enhance generalisation across phases.
+[buzzega2020dark], ER addresses the issue of catastrophic forgetting by allowing the model to retain a subset of data from earlier phases.
+During phase $t$, the model is trained not only on the current task data $\mathcal{D}^{t}$, but also on a set of exemplars $\mathcal{E}^{1:t-1}$, collected from previous phases.
+These exemplars are accumulated incrementally: $\mathcal{E}^{1:t-1}=\mathcal{E}^{1}\cup\dots\cup\mathcal{E}^{t-1}.$ Here, $\mathcal{E}^{t}$ denotes the exemplar subset stored after phase $t$, and $|\mathcal{E}^{t}|$ denotes the exemplar replay size.
+The combined replay dataset is denoted as: ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}=\mathcal{D}^{t}\cup\mathcal{E}^{1:t-1}.$ Epoch training consists of two stages: first, the model is trained on the current data, $\mathcal{D}^{t}$; then it is trained using exemplars from $\mathcal{E}^{1:t-1}$.
+While previous CI-3DIS approaches rely on large exemplar sets to mitigate forgetting [boudjoghra20243d], such strategies become impractical under memory constraints.
+Consequently, although ER aids knowledge retention, it is insufficient on its own for robust performance with a limited exemplar budget.
+Therefore, PLG and CBR are introduced to better preserve past representations and enhance generalisation across phases.
 Pseudo-Label Generator (PLG).
-In the CI-3DIS setting, at phase $t$ (for $t>1$), although we no longer have access to the ground-truth labels for previously seen classes, the model from the previous phase, $\Phi^{t-1}$, remains available and retains knowledge acquired from past tasks.
-We use this model to generate pseudo-labels for previous classes, thereby providing approximate past supervision during training.
-This allows the current model $\Phi^{t}$ to retain past knowledge and reduces forgetting, even without access to original annotations.
-Given a point cloud and label pair $(\mathbf{P},\mathbf{Y}^{t})$ from $\mathcal{D}^{t}$, the previous model model $\Phi^{t-1}$ generates pseudo-labels for the previously learned classes as $\hat{\mathbf{Y}}^{1:t-1}=\Phi^{t-1}(\mathbf{P})$, which we denote as $\tilde{\mathbf{Y}}^{t}$ for brevity.
-However, we observe that using all predictions from $\Phi^{t-1}$ introduces noisy or incorrect labels, which degrade overall performance.
+In the CI-3DIS setting, during phase $t$ (where $t>1$), although ground-truth labels for previously seen classes are unavailable, the model from the previous phase $\Phi^{t-1}$, which preserves knowledge of past tasks, is retained.
+We use this model to generate pseudo-labels for previously seen classes, thereby providing approximate supervision during training.
+This allows the current model $\Phi^{t}$, to retain prior knowledge and reduce forgetting, even without access to ground-truth annotations.
+Given a point cloud and label pair $(\mathbf{P},\mathbf{Y}^{t})$, from $\mathcal{D}^{t}$, the previous model $\Phi^{t-1}$, generates pseudo-labels for the previously seen classes as $\hat{\mathbf{Y}}^{1:t-1}=\Phi^{t-1}(\mathbf{P})$, which we denote as $\tilde{\mathbf{Y}}^{t}$ for brevity.
+However, we observe that utilising all predictions from $\Phi^{t-1}$ introduces noisy or incorrect labels, which degrade overall performance.
 To mitigate this, we select the top-$K$ most confident instance predictions, denoted as $\tilde{\mathbf{Y}}^{t}_{K}$, and concatenate them with the ground-truth labels of the current task to obtain the supervision set: ${\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{\mathbf{Y}}^{t}}=\mathbf{Y}^{t}\mathbin{\|}{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\tilde{\mathbf{Y}}^{t}_{K}}$.
-At the end, Using this supervision set, the model is optimized via the following updated objective:
-$\mathcal{L}_{\text{CI-3DIS}}({\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}};\Phi^{t})$ $=\frac{1}{|{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}}\frac{1}{|{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{\mathbf{Y}}^{t}}|}\sum_{({\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{m}_{j}^{t}},{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{c}_{j}^{t}})\in{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{\mathbf{Y}}^{t}}}$
-$\quad\left(\mathcal{L}_{\text{mask}}({\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{m}_{j}^{t}},\hat{m}_{j}^{t})+\lambda\cdot\mathcal{L}_{\text{cls}}({\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{c}_{j}^{t}},\hat{c}_{j}^{t})\right),$ (4)
-where the loss is computed over both the current task’s ground-truth labels and the top-$K$ confident pseudo-labels from previous phases, enabling the model to retain prior knowledge while adapting to new classes.
+Following this, the loss is computed over both the current task’s ground-truth labels and the top-$K$ confident pseudo-labels from earlier phases, enabling the model to retain prior knowledge while adapting to new classes.
 Class-Balanced Re-weighting (CBR).
-While ER and PLG contribute to preserving prior knowledge, we observe that the model’s predictions are biased toward frequently occurring classes, leading to forgetting of rare categories.
-This issue is further reinforced by the top-$K$ pseudo-labels selected by PLG, which predominantly represent dominant classes in the data.
-To alleviate this issue, we introduce a Class-Balanced Re-weighting (CBR) scheme that adjusts for class imbalance based on frequency of object occurrence.
-However, at task $t$, we only have access to the dataset $\mathcal{D}^{t}$ and its corresponding label distribution $p^{t}(c)$ over classes $c\in\mathcal{C}^{t}$.
-The datasets and class distributions from previous tasks $\{\mathcal{D}^{i},p^{i}(c)\}_{i=1}^{t-1}$ are no longer accessible – making it challenging to directly account for class imbalance across all seen tasks.
+While ER and PLG contribute to preserving prior knowledge, we observe that the model’s predictions are biased towards frequent classes, resulting in the forgetting of rare categories.
+This issue is further amplified by the top-$K$ pseudo-labels selected by PLG, which predominantly represent dominant classes in the data.
+To address this, we propose a Class-Balanced Re-weighting (CBR) scheme that compensates for class imbalance using object frequency statistics.
+At task $t$, only the current dataset $\mathcal{D}^{t}$ and its label distribution $p^{t}(c)$ over classes $c\in\mathcal{C}^{t}$ are available.
+Datasets and class distributions from earlier tasks $\{\mathcal{D}^{i},p^{i}(c)\}_{i=1}^{t-1}$ are no longer accessible, which makes it challenging to directly account for class imbalance across all previously seen tasks.
 Therefore, we propose to leverage the pseudo-label predictions of the frozen model $\Phi^{t-1}$ as a proxy for the class distribution across previously learned categories.
-During each training iteration, we apply the previous model $\Phi^{t-1}$ to the current input point cloud $\mathbf{P}$ to generate pseudo-labels $\tilde{\mathbf{Y}}^{t}$ , from which we accumulate class-wise frequency statistics corresponding to prior tasks.
+During each training iteration, the previous model $\Phi^{t-1}$ is applied to the current input $\mathbf{P}$ to generate pseudo-labels $\tilde{\mathbf{Y}}^{t}$, from which we accumulate class-wise frequency statistics for previously learned classes.
 Let $\mathcal{C}^{1:t}=\bigcup_{i=1}^{t}\mathcal{C}^{i}$ represent the union of all classes seen up to task $t$.
-At the end of each epoch, we compute the overall class frequency $\mathbf{f}\in\mathbb{R}^{|\mathcal{C}^{1:t}|}$ by combining the pseudo-label distribution $\tilde{p}^{t}(c)$ derived from $\tilde{\mathbf{Y}}^{t}$ with the observed ground-truth label distribution $p^{t}(c)$ from the current task.
-The resulting frequency vector $\mathbf{f}=[f(c)]_{c\in\mathcal{C}^{1:t}}$ is then used to compute class-wise $\mathbf{w}=[w(c)]_{c\in\mathcal{C}^{1:t}}$ weights for balanced pseudo-label selection and to tune current model to favour rare classes.
-During the next epoch, the model $\Phi^{t-1}$ produces soft pseudo-label predictions $\tilde{\mathbf{Y}}^{t}$ , where each component $\tilde{\mathbf{Y}}^{t}[c]$ denotes the confidence score for class $c\in\mathcal{C}^{1:t-1}$.
-To promote balanced pseudo-label selection, we apply class-wise re-weighting using $w(c)=\frac{1}{f(c)+\epsilon}$, where $\epsilon$ is a small constant for numerical stability.
-The adjusted predictions are computed by applying the class-wise weights as: $\tilde{\mathbf{Y}}^{t,w}[c]=w(c)\cdot\tilde{\mathbf{Y}}^{t}[c],\quad\forall c\in\mathcal{C}^{1:t-1}.$ We then perform top-$K$ selection over both the original and the re-weighted scores, yielding two pseudo-label sets: $\tilde{\mathbf{Y}}^{t}_{K}$ and $\tilde{\mathbf{Y}}^{t,w}_{K}$, respectively.
-The final supervision label set $\overline{\mathbf{Y}}^{t}$ is constructed by concatenating the ground-truth labels $\mathbf{Y}^{t}$ with both sets of pseudo-labels: ${\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}=\mathbf{Y}^{t}\mathbin{\|}{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\tilde{\mathbf{Y}}^{t}_{K}}\mathbin{\|}{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\tilde{\mathbf{Y}}^{t,w}_{K}}.$
+At the end of each epoch, we compute the overall class frequency $\mathbf{f}\in\mathbb{R}^{|\mathcal{C}^{1:t}|}$ by combining the pseudo-label distribution $\tilde{p}^{t}(c)$ derived from $\tilde{\mathbf{Y}}^{t}$ with the ground-truth label distribution $p^{t}(c)$ from the current task.
+The resulting frequency vector $\mathbf{f}=[f(c)]_{c\in\mathcal{C}^{1:t}}$ is then used to compute class-wise weights $\mathbf{w}=[w(c)]_{c\in\mathcal{C}^{1:t}}$, which guide balanced pseudo-label selection and help the model focus on rare classes.
+In the next epoch, the model $\Phi^{t-1}$ produces soft pseudo-label predictions $\tilde{\mathbf{Y}}^{t}$ (as described in PLG) , where each component $\tilde{\mathbf{Y}}^{t}[c]$ denotes the confidence score for class $c\in\mathcal{C}^{1:t-1}$.
+To promote balanced pseudo-label selection, class-wise re-weighting is applied using $w(c)=\frac{1}{f(c)+\epsilon}$, where $\epsilon$ is a small constant for numerical stability.
+The adjusted predictions are computed by applying the class-wise weights as: $\tilde{\mathbf{Y}}^{t,w}[c]=w(c)\cdot\tilde{\mathbf{Y}}^{t}[c],\quad\forall c\in\mathcal{C}^{1:t-1}.$ Top-$K$ selection is then applied to both the original and the re-weighted scores, yielding two pseudo-label sets: $\tilde{\mathbf{Y}}^{t}_{K}$ and $\tilde{\mathbf{Y}}^{t,w}_{K}$, respectively.
+The final augmented target set, $\overline{\mathbf{Y}}^{t}$, is constructed by concatenating the ground-truth labels $\mathbf{Y}^{t}$ with both sets of pseudo-labels:
+${\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}=\mathbf{Y}^{t}\mathbin{\|}{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\tilde{\mathbf{Y}}^{t}_{K}}\mathbin{\|}{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\tilde{\mathbf{Y}}^{t,w}_{K}}.$ (2)
 Although the above re-weighting mitigates bias from prior tasks, it does not address class imbalance within the current task $t$.
-To resolve this, we extend the class-balanced weights $w_{c}$ to incorporate statistics from both previously seen and current classes, i.e., over $\mathcal{C}^{1:t}$ rather than only $\mathcal{C}^{1:t-1}$, and denote the updated weights as $w_{c}^{\prime}$, which are then integrated into the training objective.
-After incorporating all components, the modified objective in Eq.
-1 becomes the final updated objective used in CLIMB-3D:
-$\mathcal{L}_{\text{CI-3DIS}}({\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}};\Phi^{t})$ $=\frac{1}{|{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}}\frac{1}{|{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}|}\sum_{({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{m}_{j}^{t}},{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{c}_{j}^{t}})\in{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}}$
-$\quad\left(\mathcal{L}_{\text{mask}}({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{m}_{j}^{t}},\hat{m}_{j}^{t})+{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime\prime}_{c}}\cdot\mathcal{L}_{\text{cls}}({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{c}_{j}^{t}},\hat{c}_{j}^{t})\right),$ (5)
-where ${\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime\prime}_{c}}={\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime}_{c}}\cdot\lambda_{\mathrm{cls}}$ denotes the scaled weight vector, and $\lambda_{\mathrm{cls}}$ is a loss-balancing hyperparameter.
-To address the class-imbalance across tasks, our CBR module jointly re-weights pseudo-label selection and classification loss, ensuring balanced supervision across all seen categories, including rare ones.
-This unified strategy reduces bias and enhances performance, particularly for under-represented classes.
+To resolve this, the class-balanced weights $w_{c}$ are extended to incorporate statistics from both previously seen and current classes, i.e., over $\mathcal{C}^{1:t}$ rather than only $\mathcal{C}^{1:t-1}$.
+The updated weights, denoted as $w_{c}^{\prime}$, are used to re-weights both pseudo-label selection and classification loss, ensuring balanced supervision across all seen categories, including rare ones.
+This unified strategy reduces bias and enhances performance across phases.
+The pseudo-code for constructing the supervision set in Eq.˜2, which incorporates both PLG and CBR components, is provided in Appendix A.
+Final Objective. The final training objective of CLIMB-3D, incorporating ER, PLG, and CBR, is formulated in Eq.˜1 as:
+$\mathcal{L}({\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}};\Phi^{t})=\frac{1}{|{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}}\frac{1}{|{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}|}\sum_{({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{m}_{j}^{t}},{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{c}_{j}^{t}})\in{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}}\left(\mathcal{L}_{\text{mask}}({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{m}_{j}^{t}},\hat{m}_{j}^{t})+{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime\prime}_{c}}\cdot\mathcal{L}_{\text{cls}}({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{c}_{j}^{t}},\hat{c}_{j}^{t})\right),$ (3)
+where ${\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime\prime}_{c}}={\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime}_{c}}\cdot\lambda_{\mathrm{cls}}$ denotes the scaled weight vector, and $\lambda_{\mathrm{cls}}$ is a hyperparameter for balancing the loss terms.
 3.3 Benchmarking Incremental Scenarios
-While CI-3DIS methods offer a wide range of practical applications, they frequently rely on the assumption of uniform sample distribution – an assumption that rarely holds in real-world settings.
+While CI-3DIS methods offer a wide range of practical applications, they frequently rely on the assumption of uniform sample distribution, which rarely holds in real-world settings.
 In practice, the number of object categories, denoted by $\mathcal{C}$, is often large and characterised by substantial variability in category frequency, shape, structure, and size.
 To address these challenges, we propose three incremental learning scenarios, each designed to capture a different facet of real-world complexity.
-The design of these scenarios is detailed below; for more information and illustration, please refer Appendix Appendix C of the supplementary material.
-Split_A: Frequency Scenarios.
-This scenario acknowledges that datasets are often labelled based on the frequency of category occurrences.
-To accommodate this, we propose a split where the model initially learns from the most frequent categories and subsequently incorporates the less frequent ones in later stages.
-By prioritizing the training of frequently occurring categories, the model can establish a strong foundation before expanding its knowledge to handle rare categories.
-Split_B: Semantic Scenarios.
+The design of these scenarios is detailed below; for further information and illustrations, refer to Appendix E of the supplementary material.
+Split-A: Frequency Scenarios.
+This scenario acknowledges that datasets are often labelled based on the frequency of categories.
+To accommodate this, we propose a split where the model learns from the most frequent categories and subsequently incorporates the less frequent ones in later stages.
+By prioritising the training of frequently occurring categories, the model can establish a strong foundation before expanding to handle rare categories.
+Split-B: Semantic Scenarios.
 Beyond frequency, semantic similarity is crucial in real-world deployments.
-While objects often share visual or functional traits, models may encounter semantically different categories in new environment.
-To simulate this challenge, we introduce the Split_Bscenario.
+While objects often share visual or functional traits, models may encounter semantically different categories in new environments.
+To simulate this challenge, we introduce the Split-B scenario.
 Here, categories are grouped based on their semantic relationships, and the model is incrementally trained on one semantic group at a time.
 This setup encourages the model to generalise across semantically similar categories and adapt more effectively when exposed to new ones.
-Unlike the Split_Ascenario, which organises learning based on category frequency, the Split_Bscenario may contain both frequent and infrequent categories within a single task, focusing instead on semantic continuity and transfer.
-Split_C: Random Scenarios.
+Unlike the Split-A scenario, which organises learning based on category frequency, the Split-B scenario may contain both frequent and infrequent categories within a single task, focusing on semantic continuity and transfer.
+Split-C: Random Scenarios.
 In some cases, data labelling is driven by the availability of objects rather than predefined criteria.
-To capture this, we introduce the Split_Cscenario, which represents a fully random setting where any class can appear in any task, resulting in varying levels of class imbalance.
+To capture this, we introduce the Split-C scenario, which represents a fully random setting where any class can appear in any task, resulting in varying levels of class imbalance.
 By exposing the model to such diverse and unpredictable distributions, we aim to improve its robustness in real-world situations where labelled data availability is inconsistent.
-By designing these incremental scenarios, we aim to provide a more realistic representation of object distributions, frequencies, and dynamics encountered in the real world.
+These incremental scenarios are designed to provide a more realistic representation of object distributions, frequencies, and dynamics encountered in the real world.
 4 Experiments
+4.1 Setup
+Datasets.
+We evaluate CLIMB-3D on ScanNet200 [rozenberszki2022language], which comprises 200 object categories and exhibits significant class imbalance, which makes it ideal for simulating and assessing real-world scenarios.
+In addition, we compare our approach against existing incremental learning methods using ScanNetV2 [dai2017scannet] in the 3DSS setting.
+For this evaluation, we follow the standard training and validation splits defined in prior works [Yang_2023_CVPR] to ensure consistency with existing methods.
+Evaluation Metrics.
+We evaluate our method using Mean Average Precision (mAP), a standard metric for 3DIS which provides a comprehensive measure of segmentation quality by accounting for both precision and recall.
+For comparison with existing 3DSS methods, we report the mean Intersection over Union (mIoU), which quantifies the overlap between predicted and ground truth segments.
+To assess the model’s ability to mitigate catastrophic forgetting in incremental settings, we use the Forgetting Percentage Points (FPP) metric, as defined in [liu2023continual].
+FPP captures performance degradation by measuring the accuracy drop on the initially seen categories between the first and final training phases.
+Detailed descriptions of the incremental scenarios (Split-A, Split-B, Split-C) and implementation are provided in Appendix B.
+4.2 Results and Discussion
 Table 1: Comparison between the proposed method and the baseline in the 3DIS setting, evaluated using mAP25, mAP50, mAP, and FPP after training across all phase.
@@
 Baseline 16.46 14.29 10.44 51.30 46.82
-Split_A CLIMB-3D (Ours) 35.69 31.05 22.72 3.44 2.63
+Split-A CLIMB-3D (Ours) 35.69 31.05 22.72 3.44 2.63
 Baseline 17.22 15.07 10.93 46.27 42.1
-Split_B CLIMB-3D (Ours) 35.48 31.56 23.69 8.00 5.51
+Split-B CLIMB-3D (Ours) 35.48 31.56 23.69 8.00 5.51
 Baseline 25.65 21.08 14.85 31.68 28.84
-Split_C CLIMB-3D (Ours) 31.59 26.78 18.93 9.10 7.89
+Split-C CLIMB-3D (Ours) 31.59 26.78 18.93 9.10 7.89
+To evaluate our proposed approach, we construct a baseline using ER [buzzega2020dark] for the 3DIS setting, as no existing incremental baselines are available for this task.
+In contrast, for the 3DSS setting, where incremental baselines do exist, we compare our method against three prior approaches [serra2018overcoming, Yang_2023_CVPR, li2017learning].
+Results on CI-3DIS.
+Tab.˜1 compares CLIMB-3D against the ER baseline across the three CI-3DIS scenarios, evaluated after all phases using $\text{mAP}_{25}$, $\text{mAP}_{50}$, overall mAP, and FPP.
+In the Split-A scenario, characterised by significant distribution shifts, our method achieves gains of +19.23%, +16.76%, and +12.28% in $\text{mAP}_{25}$, $\text{mAP}_{50}$, and overall mAP, respectively, while drastically reducing forgetting with FPP scores of 3.44% ($\text{mAP}_{25}$) and 2.63% ($\text{mAP}_{50}$), both over 45 points lower than the baseline.
+Under the Split-B scenario, with semantically related new categories, improvements of +18.26%, +16.49%, and +12.76% are observed alongside a reduction in forgetting to 8.00% and 5.51% from baseline levels above 42%.
+For the Split-C scenario, featuring increasing geometric complexity, CLIMB-3D maintains solid gains of +5.94%, +5.70%, and +4.08% across mAP metrics, with forgetting reduced by over 22 points.
+Overall, these results demonstrate robust forward transfer, minimal forgetting, and stable learning, highlighting the method’s effectiveness for scalable continual 3D semantic understanding.
+Results on CI-3DSS.
 Table 2: Comparison between the proposed method and existing baselines in the 3DSS setting on ScanNetV2 [dai2017scannet], evaluated using mIoU.
@@
 LwF [li2017learning] 30.38 13.37 26.13
-Yang Yang_2023_CVPR 34.16 13.43 28.98
+[…]
 CLIMB-3D (Ours) 69.39 32.56 59.38
-Table 3: Ablation study showing the impact of each component ER, PLG, and CBR in a three-phase setup.
-Each split (s) and number corresponds to data introduced at that phase.
-The All column reports performance over all classes seen so far.
-Joint training (Oracle) is highlighted in gray, and the full model in orange.
-Best results are in bold.
+Although our primary focus is on CI-3DIS, we also evaluate our method under the CI-3DSS setting to enable comparisons with existing approaches.
+To do so, we adapt our predictions to assign each point the label corresponding to the highest-confidence mask and exclude background classes (e.g., floor and wall), as these are not part of the semantic segmentation targets.
+Tab.˜2 presents a comparison between our method and existing baselines on the ScanNet V2 dataset [dai2017scannet], evaluated using mIoU across two training phases and overall.
+Although originally designed for instance segmentation, our method generalises effectively to semantic segmentation, achieving substantial gains of +35.23% mIoU in Phase 1 and +19.1% in Phase 2.
+Overall, it achieves 59.38% mIoU, significantly higher than the $\sim$30% mIoU of prior methods, highlighting its robustness and transferability.
+Refer to Appendix C, D, and F of the supplementary material for detailed analyses of per-phase performance, rare-class evaluation, and qualitative result comparisons, respectively.
+4.3 Ablation
+Table 3: Ablation study illustrating the impact of each component in a three-phase setup.
+Each split (s) and corresponding number indicates data introduced at that phase.
+Best results are highlighted in bold.
 Row Modules p=1 $\uparrow$ p=2 $\uparrow$ p=3 $\uparrow$ FPP $\downarrow$
@@
 5. + CBR 56.82 54.67 33.75 44.13 54.19 12.02 26.55 31.05 2.63
-4.1 Setup
-Datasets.
-We evaluate our method on the ScanNet200 [rozenberszki2022language], which comprises 200 object categories and exhibits significant class imbalance – making it ideal for simulating and assessing real-world scenarios.
-In addition, we benchmark our approach against existing incremental learning methods using the ScanNetV2 [dai2017scannet] in a 3DSS setting.
-For this evaluation, we follow the standard training and validation splits defined in prior works [Yang_2023_CVPR].
-Evaluation Metrics.
-We evaluate our method using mean Average Precision (mAP), a standard metric for 3DIS that provides a comprehensive measure of segmentation quality by accounting for both precision and recall.
-For comparison with existing 3DSS methods, we report the mean Intersection over Union (mIoU), which quantifies the overlap between predicted and ground truth segments.
-To assess the model’s ability to mitigate catastrophic forgetting in incremental settings, we use the Forgetting Percentage Points (FPP) metric, as defined in [liu2023continual].
-FPP captures performance degradation by measuring the accuracy drop on the initially seen categories between the first and final training phases.
-Incremental Scenarios.
-As discussed in Section 3.3, we design three incremental scenarios – Split_A, Split_B, and Split_C– each with three tasks, based on object frequency, semantic similarity, and random grouping, respectively.
-For Split_A, we follow the class split provided by rozenberszki2022language, with a class distribution of 66-68-66 across the three tasks.
-In Split_B, we partition the classes into 74-50-76 based on semantic similarity, which is calculated using the CLIP [radford2021learning] text encoder, followed by clustering using K-Means.
-Finally, in Split_C, classes are randomly shuffled and split into three sets, with 67-67-66 categories per split.
-These scenarios comprehensively evaluate our approach’s performance and generalization in varied real-world conditions.
-Implementation Details.
-For all the experiments we use Mask3D [Schult23ICRA] with the Minkowski Res16UNet34C [choy20194d] as a feature backbone.
-We adopt the data augmentation, hyperparameters, and training strategy described in [Schult23ICRA], with using the AdamW optimizer [loshchilov2017decoupled] and a one-cycle learning rate scheduler [smith2018superconvergence], and results are evaluated on the entire validation set.
-In incremental training, learning is divided into three phases, introducing one split per phase across the three scenarios described in Sec.
-3.3.
-For each phase, we train for 500, 300, and 300 epochs, respectively, except for the Oracle setting (Row 1 in Tab.
-3), which is trained once for 600 epochs.
-We retain the same hyperparameters throughout, modifying only the number of training epochs.
-The exemplar replay size is fixed at $|\mathcal{E}^{t}|=50$ scenes.
-4.2 Results and Discussion
-To evaluate our proposed approach, we construct a baseline using ER [buzzega2020dark] for the 3DIS setting, as no existing incremental baselines are available for this task.
-In contrast, for the 3DSS setting, where prior incremental baselines do exist, we compare our method against three existing approaches [serra2018overcoming, Yang_2023_CVPR, li2017learning].
-Results on CI-3DIS.
-Tab.
-1 compares CLIMB-3D against the ER baseline across three CI-3DIS scenarios, evaluated after all phases using $\text{mAP}_{25}$, $\text{mAP}_{50}$, overall mAP, and FPP.
-In the Split_Ascenario, characterized by significant distribution shifts, it achieves gains of +19.23%, +16.76%, and +12.28% in $\text{mAP}_{25}$, $\text{mAP}_{50}$, and overall mAP, while drastically reducing forgetting with FPP scores of 3.44% and 2.63%, over 47 points lower than the baseline.
-Under the Split_Bscenario, with semantically related new categories, improvements of +18.26%, +16.49%, and +12.76% are observed alongside a reduction in forgetting to 8.00% and 5.51% from baseline levels above 42%.
-For the Split_Cscenario, featuring increasing geometric complexity, it maintains solid gains of +5.94%, +5.70%, and +4.08% across mAP metrics, with forgetting reduced by over 22 points.
-Overall, these results demonstrate robust forward transfer, minimal forgetting, and stable learning, validating the method’s effectiveness for scalable continual 3D semantic understanding.
-Results on CI-3DSS.
-Although our primary focus is on CI-3DIS, we additionally evaluate our method under the CI-3DSS setting to enable comparison with existing approaches.
-For this, we adapt our predictions by assigning each point the label corresponding to the highest-confidence mask and exclude background classes (floor and wall), as these are not part of the semantic segmentation targets.
-Tab.
-2 presents a comparison between our method and existing baselines on the ScanNet V2 dataset [dai2017scannet], evaluated using mIoU across two training phases and overall.
-Despite being designed for instance segmentation, our method generalises effectively to semantic segmentation, achieving substantial gains: +35.23% mIoU in Phase 1 and +19.1% in Phase 2.
-Overall, it achieves 59.38% mIoU – significantly higher than the $\sim$30mIoU of prior methods – highlighting its robustness and transferability.
-For a detailed per-phase analysis, evaluation of rare classes, and qualitative result comparisons, please refer to Appendix Appendix A, Appendix Appendix B, and Appendix Appendix D, respectively, in the supplementary material.
-4.3 Ablation
 We conduct an ablation study to evaluate the contribution of each component in our framework.
-As an upper bound, we report the performance of an Oracle model trained jointly on the full dataset.
+As an upper bound, we report the performance of an Oracle model, which is trained jointly on the full dataset.
 For the incremental setting, we follow the previously defined splits and first train the model naively across phases.
 We then incrementally add each module to isolate its impact.
-Tab.
-3 presents results on the Split_Ascenario using both $\text{mAP}_{50}$  and FPP metrics.
+Tab.˜3 presents results for the Split-A scenario using both $\text{mAP}_{50}$  and FPP metrics.
 Naïve Training.
 When trained naïvely without any dedicated modules, the model suffers from severe catastrophic forgetting, as evident in row 2.
-The model completely forgets previously learned classes upon entering new phases, with performance dropping to zero for earlier splits and a high FPP of 56.82.
+The model entirely forgets previously learned classes upon entering new phases, with performance dropping to zero for earlier splits and FPP rising to 56.82.
 Effect of ER.
@@
 Effect of PLG.
-The addition of the pseudo-label generator (PLG), which generates labels for previous classes by retaining a copy of the model from earlier phase, facilitates knowledge retention and forward transfer.
+The addition of the pseudo-label generator (PLG), which generates labels for previous classes by retaining a copy of the model from an earlier phase, facilitates knowledge retention and forward transfer.
 As shown in row 4, PLG significantly reduces forgetting and enhances performance on current tasks.
@@
 Effect of CBR.
-Finally, class-balanced re-weighting (CBR) mitigates class imbalance both during pseudo-labelling and current task learning by adjusting each class’s contribution based on its frequency.
-As shown in row 5, CBR enhances s1 retention over PLG in phases 2 and 3 (+4.67% and +4.41%) and further improves performance on s2 and s3.
-It also achieves the lowest FPP of 2.63, reflecting effective forgetting mitigation.
-Overall, CBR offers the best trade-off between retaining past knowledge and learning new tasks, outperforming PLG and ER by 4.41% and 43.81%, respectively.
+Finally, class-balanced re-weighting (CBR) mitigates class imbalance during both pseudo-labelling and current task learning by adjusting each class’s contribution based on its frequency.
+As shown in row 5, CBR enhances s1 retention over PLG in Phases 2 and 3 (+4.67% and +4.41%), and further improves performance on s2 and s3.
+It also achieves the lowest FPP of 2.63, reflecting strong forgetting mitigation.
+Overall, CBR offers the best trade-off between retaining past knowledge and acquiring new tasks, outperforming PLG and ER by 4.41% and 43.81%, respectively.
+4.4 Limitations & Future Work
+While CLIMB-3D achieves strong performance across CI-3DIS and CI-3DSS settings, several limitations remain and warrant discussion.
+The experiments are currently limited to indoor datasets (ScanNet200 and ScanNetV2), therefore evaluating performance on outdoor scenes would help assess generalisability across diverse environments.
+The current setup also considers only three incremental phases, whereas real-world applications often involve longer sequences; incorporating such extended setups would better reflect practical challenges.
+Moreover, the model operates in a uni-modal setting.
+Integrating multi-modal cues, such as vision-language models, could further enhance performance, as suggested by recent 2D and 3D studies [hong20233d, thengane2022clip, thengane2025foundational].
 5 Conclusion
-We address the critical challenge of catastrophic forgetting in class-incremental 3D instance segmentation by proposing a modular framework that combines exemplar replay (ER), a pseudo-label generator (PLG), and class-balanced re-weighting (CBR).
-Our method incrementally adapts to new classes while preserving knowledge from previous tasks without requiring access to the full dataset.
-Through extensive experiments on a three-phase benchmark, we demonstrate the individual and combined effectiveness of each component.
-ER provides a easy knowledge storage, PLG enhances knowledge retention via pseudo-label supervision, and CBR further refines learning by mitigating class imbalance.
-Together, these modules significantly reduce forgetting and improve overall segmentation performance across all phases.
-By balancing stability and plasticity, our method contributes to continual 3D scene understanding and establishing a new baseline for future research.
+This work addresses the challenge of catastrophic forgetting in class-incremental 3D instance segmentation by introducing a modular framework that integrates exemplar replay (ER), a pseudo-label generator (PLG), and class-balanced re-weighting (CBR).
+The proposed method incrementally adapts to new classes while preserving prior knowledge, all without requiring access to the full dataset.
+Extensive experiments on a three-phase benchmark validate the individual and combined effectiveness of each component: ER enables efficient memory-based retention; PLG facilitates knowledge preservation through pseudo-supervision; and CBR mitigates class imbalance to enhance learning stability.
+Together, these components significantly reduce forgetting and improve segmentation performance across all phases.
+By achieving a strong balance between stability and plasticity, our approach advances continual 3D scene understanding and sets a new baseline for future research in this area.
 Supplementary Material
-Table 4: Comparison of results in terms of mAP50 with and proposed CLIMB-3D for three different scenarios.
-Each scenario is trained in three phases ($\mathtt{phase}=1,2,3$) by introducing a single split s at a time.
-The results highlighted in orange are with the proposed method, and the best results for each scenario are in bold.
+Appendix Appendix A Illustrating Incremental scenarios
+Figure 3: Tasks are grouped into incremental scenarios based on object frequency, semantic similarity, and random assignment.
+,  , and   denote different tasks; shapes indicate object categories;   marks the background.
+Left: Grouped by category frequency.
+Middle: Grouped by semantic similarity (e.g., similar shapes).
+Right: Randomly grouped, mixing semantic and frequency variations.
+Appendix Appendix B Per Phase Analysis
+Table 4: Comparison of results in terms of mAP50 with the proposed CLIMB-3D across three different scenarios.
+Each scenario is trained over three phases ($\mathtt{phase}=1,2,3$), introducing a single split s at a time.
+Results highlighted in orange correspond to the proposed method, and the best results for each scenario are shown in bold.
 Scenarios Methods phase=1 phase=2 phase=3
@@
 Baseline 56.82 18.51 32.81 25.72 10.38 9.43 24.27 14.28
-Split_A CLIMB-3D 56.82 54.67 33.75 44.13 54.19 12.02 26.55 31.05
+Split-A CLIMB-3D 56.82 54.67 33.75 44.13 54.19 12.02 26.55 31.05
 Baseline 51.57 13.32 42.21 24.53 9.55 12.45 26.78 15.07
-Split_B CLIMB-3D 51.57 46.74 37.45 43.13 46.06 15.95 26.68 31.56
+Split-B CLIMB-3D 51.57 46.74 37.45 43.13 46.06 15.95 26.68 31.56
 Baseline 36.40 7.74 37.62 22.32 7.55 15.96 40.41 21.08
-Split_C CLIMB-3D 36.40 32.63 33.38 33.00 28.51 17.11 34.64 26.78
-Table 5: Results for classes observed by the model 1–20 times during an epoch, evaluated on Split_A  for Phase 2, in terms of $\text{mAP}_{50}$.
-Classes Seen Count ER+KD ER+KD+IC
+Split-C CLIMB-3D 36.40 32.63 33.38 33.00 28.51 17.11 34.64 26.78
+We extend the analysis from Tab.˜1 to Tab.˜4 to highlight the impact of our proposed method on individual splits across various scenarios.
+The results clearly demonstrate that our model consistently retains knowledge of previous tasks better than the baseline.
+For Split-A, our model shows improvement throughout the phase.
+In Phase 3 of (s2), although both the baseline and our method exhibit a performance drop, our method reduces forgetting significantly compared to the baseline.
+The Split-B scenario, while more complex than Split-A, achieves comparable results due to semantic similarity among classes within the same task.
+In Phase 2, our model achieves overall all 43.13% $\text{mAP}_{50}$ compared to 24.53% on baseline, a similar trend is observed in Phase 3, where our method not only consistently improves learning but also enhances retention of previous information.
+After all three tasks, our method achieves an overall performance of 31.56% AP50, compared to 15.07% for the baseline.
+In the Split-C scenario, the first-stage model struggles due to the increased complexity introduced by random grouping.
+In Phase 2, while the baseline focuses on learning the current task, it suffers from severe forgetting of prior knowledge.
+Conversely, our method balances new task learning with the retention of earlier information.
+By Phase 3, the model effectively consolidates s1 and maintains strong performance across all task splits.
+Overall, our proposed method improves mAP by 5.6%.
+In this supplementary material, we first demonstrate the performance gains on rare classes achieved by incorporating the IC module in Appendix C.
+Next, we provide detailed split information for all scenarios, based on class names, in Appendix D.
+Finally, we present a qualitative comparison between the baseline method and our proposed approach in Appendix E.
+Appendix Appendix C Evaluation on Rare Categories
+The proposed imbalance correction (IC) module, as detailed in Section 4.2, is designed to address the performance gap for rare classes.
+To assess its impact, we compare its performance with the framework which has exemplar replay (ER) and knowledge distillation (KD).
+Specifically, we focus on its ability to improve performance for rare classes, which the model encounters infrequently compared to more common classes.
+Table 5: Results for classes observed by the model 1–20 times during an epoch, evaluated on Split-A  for Phase 2, in terms of $\text{mAP}_{50}$.
+Classes Seen Count ER + PLG ER + PLG + CBR
 paper towel dispenser 2 73.10 74.90
@@
 Average - 46.66 54.98
-Table 6: Results for classes observed by the model 1–10 times during an epoch, evaluated on Split_A  for Phase 3, in terms of $\text{mAP}_{50}$.
+The results, shown in Tab.˜5 and Tab.˜6, correspond to evaluations on Split-A  for Phase 2 and Phase 3, respectively.
+In Phase 2, we evaluate classes seen 1–20 times per epoch, while Phase 3 targets even less frequent classes, with observations limited to 1–10 times per epoch.
+As illustrated in Tab.˜5, the IC module substantially improves performance on rare classes in terms of $\text{mAP}_{50}$  in Phase 2 of Split-A.
+For instance, classes like recycling bin and trash bin, seen only 3 and 7 times, respectively, shows significant improvement when the IC module is applied.
+Overall, the IC module provides an average boost of 8.32%, highlighting its effectiveness in mitigating class imbalance.
+Table 6: Results for classes observed by the model 1–10 times during an epoch, evaluated on Split-A  for Phase 3, in terms of $\text{mAP}_{50}$.
 Classes Seen Count ER+KD ER+KD+IC
@@
 Average - 32.08 44.21
-Figure 3: Tasks are grouped into incremental scenarios based on object frequency, semantic similarity, and random assignment.
-,  , and   denote different tasks; shapes indicate object categories;   marks the background.
-Left: Grouped by category frequency.
-Middle: Grouped by semantic similarity (e.g., similar shapes).
-Right: Randomly grouped, mixing semantic and frequency variations.
+Similarly, Tab.˜6 presents results for Phase 3, demonstrating significant gains for infrequent classes.
+For example, even though the classes such as piano, bucket, and laundry basket are observed only once, IC module improves the performance by 52.30%, 10.40%, and 13.60%, respectively.
+The ER+KD module does not focus on rare classes like shower and toaster which results in low performance, but the IC module compensates for this imbalance by focusing on underrepresented categories.
+On average, the addition of the proposed IC module into the framework outperforms ER+KD by 12.13%.
 Table 7: Classes grouped by tasks for each proposed scenario on the ScanNet200 dataset labels.
-The three scenarios Split_A, Split_A, and Split_C  are each divided into three tasks: Task 1, Task 2, and Task 3.
+The three scenarios Split-A, Split-A, and Split-C  are each divided into three tasks: Task 1, Task 2, and Task 3.
 Split_A Split_B Split_C
@@
 potted plant
-In this supplementary material, we first present the per-phase performance of the proposed CLIMB-3D model in Appendix Appendix A, followed by the analysis of gains on rare classes achieved through the incorporation of the CBR module in Appendix Appendix B.
-Next, we provide detailed split information for all scenarios, based on class names, in Appendix Appendix C.
-Finally, we present a qualitative comparison between the baseline method and our proposed approach in Appendix Appendix D.
-Appendix Appendix A Per-Phase Performance Analysis
-We extend the analysis presented in Tab.
-1 (Sec.
-4.2 of the main paper) to Tab.
-4 to examine the impact of our proposed method across individual splits and phases in different scenarios.
-The results consistently show that our method outperforms the baseline by better retaining knowledge from earlier tasks.
-Split_A.
-In the Split_Ascenario, our model shows notable improvements across all phases.
-Particularly in Phase 3, although both methods show a drop in performance for s2, our approach significantly reduces forgetting, maintaining a higher $\text{mAP}_{50}$  across all splits.
-Split_B.
-The Split_Bscenario, while more challenging due to increased semantic overlap, demonstrates our model’s strength in generalisation.
-In Phase 2, it achieves an overall $\text{mAP}_{50}$  of 43.13% compared to the baseline’s 24.53%.
-This trend continues in Phase 3, with our model achieving 31.56%, more than double the baseline’s 15.07%, indicating both improved learning and retention.
-Split_C.
-For the Split_Cscenario, where splits are randomly grouped, Phase 1 performance is initially lower.
-However, in subsequent phases, our method balances stability and plasticity better than the baseline.
-By Phase 3, it consolidates performance across all splits, improving the overall $\text{mAP}_{50}$to 26.78%, a 5.7% gain over the baseline.
-These results demonstrate the robustness of our method in handling continual learning under varied and challenging split strategies.
-Appendix Appendix B Evaluation on Rare Categories
-The proposed Class-Balanced Re-weighting (CBR) module, as detailed in 3.2, is designed to address the performance gap for rare classes.
-To assess its impact, we compare its performance with the framework which has exemplar replay (ER) and Pseudo-Label Generator (PLG).
-Specifically, we focus on its ability to improve performance for rare classes, which the model encounters infrequently compared to more common classes.
-The results, shown in Tab.
-5 and Table 6, correspond to evaluations on Split_A  for Phase 2 and Phase 3, respectively.
-In Phase 2, we evaluate classes seen 1–20 times per epoch, while Phase 3 targets even less frequent classes, with observations limited to 1–10 times per epoch.
-As illustrated in Table 5, the CBR module substantially improves performance on rare classes in terms of $\text{mAP}_{50}$  in Phase 2 of Split_A.
-For instance, classes like recycling bin and trash bin, seen only 3 and 7 times, respectively, show significant improvement when the CBR module is applied.
-Overall, the CBR module provides an average boost of 8.32%, highlighting its effectiveness in mitigating class imbalance.
-Similarly, Table 6 presents results for Phase 3, demonstrating significant gains for infrequent classes.
-For example, even though classes such as piano, bucket, and laundry basket are observed only once, the CBR module improves the performance by 52.30%, 10.40%, and 13.60%, respectively.
-The ER+PLG module does not focus on rare classes like shower and toaster, which results in low performance, but the CBR module compensates for this imbalance by focusing on underrepresented categories.
-On average, the addition of the proposed CBR module into the framework outperforms ER+PLG by 12.13%.
-Appendix Appendix C Incremental Scenarios Phases
-Table 7 presents the task splits for each proposed scenario introduced in Sec.
-3.3, and Fig.
-3 illustrates them visually for the ScanNet200 [rozenberszki2022language] dataset.
-The three scenarios – Split_A, Split_B, and Split_C– are each divided into three tasks: Task1, Task2, and Task3.
-Notably, the order of classes within these tasks is random.
-Appendix Appendix D Qualitative Results
+Appendix Appendix D Incremental Scenarios Phases
+Tab.˜7 presents the task splits for each proposed scenario introduced in Section 4.3 using the ScanNet200 dataset.
+The three scenarios, Split-A, Split-B, and Split-C, are each divided into three tasks: Task 1, Task 2, and Task 3.
+Notably, the order of classes in these tasks is random.
+Appendix Appendix E Qualitative Results
 In this section, we present a qualitative comparison of the proposed framework with the baseline method.
-Figure 4 illustrates the results on the Split_A evaluation after learning all tasks, comparing the performance of the baseline method and our proposed approach.
+Fig.˜4 illustrates the results on the Split-A evaluation after learning all tasks, comparing the performance of the baseline method and our proposed approach.
 As shown in the figure, our method demonstrates superior instance segmentation performance compared to the baseline.
@@
 Overall, our framework consistently outperforms the baseline, with several missed instances by the baseline highlighted in red circles.
-Figure 4: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split_A  evaluation after learning all tasks.
-In Figure 5, we present the results on Split_B, highlighting instances where the baseline method underperforms, marked with red circles.
+Figure 4: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split-A  evaluation after learning all tasks.
+In Fig.˜5, we present the results on Split-B, highlighting instances where the baseline method underperforms, marked with red circles.
 For example, in row 2, the baseline method incorrectly identifies the same sofa as separate instances.
@@
 In contrast, the proposed method delivers results that closely align with the ground truth, demonstrating its superior performance
-Figure 5: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split_B  evaluation after learning all tasks.
-Similarly, Figure 6 highlights the results on Split_C, where classes are encountered in random order.
+Figure 5: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split-B  evaluation after learning all tasks.
+Similarly, Fig.˜6 highlights the results on Split-C, where classes are encountered in random order.
 The comparison emphasizes the advantages of our method, as highlighted by red circles.
@@
 In contrast, our approach consistently produces results that are closely aligned with the ground truth, further underscoring its effectiveness.
-Figure 6: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split_C  evaluation after learning all tasks.
+Figure 6: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split-C  evaluation after learning all tasks.